### v1

# Towards Automated Semantic Interpretability in Reinforcement Learning via Vision-Language Models

Abstract
Semantic Interpretability in Reinforcement Learning (RL) enables transparency, accountability, and safer deployment by making the agent’s decisions understandable and verifiable. Achieving this, however, requires a feature space composed of human-understandable concepts, which traditionally rely on human specification and fail to generalize to unseen environments. In this work, we introduce Semantically Interpretable Reinforcement Learning with Vision-Language Models Empowered Automation (SILVA), an automated framework that leverages pre-trained vision-language models (VLM) for semantic feature extraction and interpretable tree-based models for policy optimization. SILVA first queries a VLM to identify relevant semantic features for an unseen environment, then extracts these features from the environment. Finally, it trains an Interpretable Control Tree via RL, mapping the extracted features to actions in a transparent and interpretable manner. To address the computational inefficiency of extracting features directly with VLMs, we develop a feature extraction pipeline that generates a dataset for training a lightweight convolutional network, which is subsequently used during RL. By leveraging VLMs to automate tree-based RL, SILVA removes the reliance on human annotation previously required by interpretable models while also overcoming the inability of VLMs alone to generate valid robot policies, enabling semantically interpretable reinforcement learning without human-in-the-loop.
I Introduction
Reinforcement learning (RL) agents operate in dynamic environments, often developing policies that rely on latent representations not directly interpretable by humans [1]. Ensuring that these learned representations align with human-understandable concepts, a goal often referred to as semantic interpretability, is crucial for ensuring safety and preventing unintended harmful actions. While progress has been made in developing interpretable AI models, achieving semantic interpretability in RL remains particularly challenging  [2].
Semantic interpretability refers to the extent to which a model’s features, decisions, or internal representations can be understood and explained in terms of human-understandable concepts. Achieving this requires two essential components: (1) the inputs to the decision-making model must be intelligible to facilitate downstream interpretation, and (2) the decision-making process must be built with an interpretable model that allows humans to understand and explain its rationale  [3]. In essence, achieving semantic interpretability involves mapping high-dimensional input data to human-understandable features that can serve as a low-dimensional input space for interpretable decision-making models.
To learn an interpretable model on a semantically interpretable feature space, various model classes can be considered, including decision trees and rule-based policies, both of which facilitate explicit reasoning about an agent’s actions. Decision trees have been used to represent policies in a structured, human-readable format [4, 5, 6]. Similarly, rule-based methods, such as logical programs and symbolic reasoning approaches, offer an alternative by encoding policies in a form that can be explicitly followed and verified [7, 8]. Additionally, hierarchical decision-making frameworks have been explored, where high-level policies generate interpretable subgoals while low-level controllers handle execution [9, 10]. However, none of these prior works address high-dimensional image or video inputs; instead, they rely either on human-defined semantic concepts or ground-truth feature vectors provided by the environment.
Existing approaches have attempted to bridge the gap between raw image inputs and interpretable features [11] to facilitate model understanding but fall short in semantic interpretability. Many approaches either produce mathematical or visual representations that do not correspond to human-understandable concepts [12, 13] or rely on predefined feature spaces that aren’t generalize across different environments, using CNNs [14] or ViTs [15] to establish mappings between inputs and semantic features. As a result, human intervention has traditionally played a key role, either by manually defining semantically meaningful feature spaces [16] or by labeling machine-generated features to associate them with semantically meaningful concepts [17].
The emergence of Vision-Language Models (VLMs) has opened new possibilities for overcoming these challenges. These models, known for their capabilities in fine-grained visual understanding, language reasoning, and decision-making  [18, 19, 20], have seen widespread use in various tasks  [19, 21]. In this work, we demonstrate that VLMs can substitute humans in defining and extracting semantic features, serving not merely as a “scarecrow in oz”  [22] or a cost-effective fallback to human labor, but as a scalable and automated solution that matches human performance.
We propose Semantically Interpretable Reinforcement Learning with Vision-Language Models Empowered Automation (SILVA), which is to the best of our knowledge the first automated framework that achieves semantic interpretability in RL and generalizes effectively to unseen environments. By leveraging VLMs to extract features from raw environment inputs and training with Interpretable Control Trees, we manage to provide the rationale behind a RL algorithm in terms of human-understandable concepts. Notably, our framework (i) maintains competitive performance with black-box models while surpassing uninterpretable models built on the same feature space, (ii) supports both image and video inputs, effectively capturing temporal information, (iii) generalizes to unseen environments without human-in-the-loop, and (iv) enables computational efficiency by training a lightweight convolutional network as an alternative to VLM-based feature extraction.
II Related Work
II-A Interpretable Reinforcement Learning
Figure 1: Overview of SILVA, our proposed framework. A. Feature Extraction involves a three-step process where: (1) The VLM identifies key features given domain knowledge and environment observation. (2) The VLM generates code to extract features from a single image frame and assesses whether certain key features require multiple frames. If so, it builds upon the single-frame code to extract these features. (3) Input-output pairs are collected using the extractor code and are used to train an Extractor Convolutional Neural Network (Extractor CNN) as a computationally efficient substitute. B. Tree-based RL Training leverages the Extractor CNN to obtain semantic features from image observations, which are subsequently used as inputs to the policy network (an Interpretable Control Tree) to generate control actions for the RL agent.
Interpretable RL is inherently challenging compared to standard supervised learning, as it involves long-term consequences to actions and latent receptions of data through exploration.  [2] A significant line of research has explored interpretability in RL through generating visual saliency maps using Jacobian-based, perturbation-based, and attribution-based methods [23, 13, 24]. However, these approaches primarily focus on post-hoc explanations and fall out of a stricter definition of Interpretable RL  [3], which requires all components including the input space and the decision-making process to be intelligible. Approaches to achieve intelligible decision-making includes interpretable tree-structures, rule and decision lists, and other rule-based methods [3], among which decision trees are arguably the easiest to verify  [25]. Among state-of-the-art tree methods, Differentiable Decision Trees (DDTs) employ a Decision Tree structure that allows for online learning via stochastic gradient descent [25], while Interpretable Continuous Control Trees (ICCTs) extend this capability to continuous control tasks, ensuring both high performance and interpretability [26]. Similarly, POETREE (Policy Extraction through Decision Trees) [27] introduces dynamically growing probabilistic decision trees that adapt throughout training, allowing the tree complexity to evolve based on the task requirements. Our work addresses the limitation of prior tree-based methods by automating feature definition and extraction, thus enabling generalization to unseen environments.
II-B Feature Extraction for Reinforcement Learning
Common techniques that map high-dimensional inputs into low-dimensional feature spaces for RL includes linear nonparametric methods, such as manifold learning and spectral learning, and nonlinear parametric methods, such as deep learning  [12]. Other work focuses on the extraction of visual features that are not inherently connected with semantic meanings.  [23] employs a self-supervised interpretable network (SSINet) to produce fine-grained attention masks for highlighting task-relevant information.  [13] computed perturbation-based saliency of the image input to deduce regions that had the most effect. However, only a limited number of studies have proposed approaches without human-in-the-loop to construct semantic feature spaces for RL training.  [28] leverages information from Atari instruction manuals to construct the feature space, then uses unsupervised object detection and grounding to acquire the corresponding
objects in the environment. However, they observe that current techniques exhibit a lack of reliability  [28]. Our framework distinguishes from prior work by enabling automatic feature definition and extraction without human-in-the-loop.
II-C VLMs for Reinforcement Learning
VLMs have been applied in various aspects of RL training, including reward design, preference feedback, and representation deriving [29, 30, 20]. One prominent approach involves using VLMs to generate reward functions.  [29] use VLMs as zero-shot reward models to specify tasks via natural language, while  [31] leverages VLMs to produce dense reward functions through code generation. Additionally,  [30, 32] use VLMs to provide preference feedback with which they updated the reward functions for preference-based RL. Another line of work leverages VLMs as feature extractors to obtain valuable information which facilitates training.  [20] extract promptable representations from VLMs and decode output tokens into embeddings, which are used to initialize RL policies. However, those representations are not derived for interpretability purposes and might not often be correct. To the best of our knowledge, our work is the first to utilize VLM for creating a semantic feature space for RL training.
III Methodology
We introduce Semantically Interpretable Reinforcement Learning with Vision-Language Models Empowered Automation (SILVA), a framework designed to automatically learn an interpretable decision-making RL policy from an unseen environment. SILVA leverages the visual comprehension and reasoning capabilities of VLMs automate feature extraction through a three-step pipeline: (1) it employs zero-shot prompting on the environment to identify a set of semantic features that deterministically contribute to the decision-making, (2) it generates feature extraction codes that captures both static and temporal information, and (3) it collects a dataset with VLM-generated extraction codes to train a lightweight convolutional neural network (Extractor CNN) as a computationally efficient substitute. Subsequently, utilizing the Extractor CNN, SILVA performs tree-based RL training with an Interpretable Control Tree, establishing a transparent mapping between the semantic feature space and action outputs to derive a semantically interpretable policy for decision-making.
III-A Feature Extraction
The feature extraction process of SILVA involves a three-step pipeline, as shown in Fig. 1. In Step 1, we query the VLM to generate key features that contribute to the decision in the environment. In Step 2, we query the VLM to generate code that extracts the identified features from the environment, either from a single image or a list of images. In Step 3, we collect a dataset from the environment with the VLM-generated extractor code, with which we train an Extractor CNN that will be used in RL training to enable batch computing and increase program speed. The methodology is formalized below.
Step 1: Querying the VLM for Semantic Features. To extract intelligible features from an unseen environment that could be used to train an effective RL policy, we need to first determine what features are useful to model decision-making. In this step, we provide the VLM with:
  • $\mathcal{D}$: Basic domain knowledge about the environment.
  • $\mathcal{I}_{0}$: An example image of the environment.
The VLM identifies key semantic features critical for decision-making, represented as two distinct sets:
  • $\mathcal{F}_{s}=\{f_{1},f_{2},\dots,f_{p}\}$: Features that can derived from a single image frame.
  • $\mathcal{F}_{m}=\{f_{p+1},f_{p+2},\dots,f_{n}\}$: Features requiring multiple frames (e.g., velocity, acceleration, tendency of change). Note that $\mathcal{F}_{m}$ may not be needed in every environment.
The process can be formalized in Eq. 1:
$\mathcal{F}=\mathcal{F}_{s}\cup\mathcal{F}_{m},\quad\mathcal{F}=\mathcal{M}(\mathcal{D},\mathcal{I}_{0}).$ (1)
For example, in the Skiing environment, $\mathcal{F}_{s}$ includes features such as the position of skier, the position of next gate, the distance to next gate, the position of nearest tree, and the distance to nearest tree, while $\mathcal{F}_{m}$ includes the current horizontal speed of the skiier. This example is also included in Fig. 1.
Step 2: Querying the VLM for Feature Extraction Code. In this step, we query the VLM to generate feature extraction code, denoted by $\mathcal{G}$, to retrieve the identified set of features $\mathcal{F}$. To mitigate hallucination, we provide visual segmentations for each image, which will be covered in the discussion section. The visual segmentation $\mathcal{S}$ is computed using a segmentation function, represented as $Seg$, such that $\mathcal{S}=\text{Seg}(\mathcal{I})$. The code generation process involves two stages:
Single Image Extraction: For features in $\mathcal{F}$, the VLM generates code $\mathcal{G}_{s}$ to extract as many features as possible from a single image-segmentation pair $(\mathcal{I},\mathcal{S})$. In practice, as demonstrated in our experiments, this approach effectively recovers all features in $\mathcal{F}_{s}$. To enhance the robustness of the generated code, the VLM is provided with multiple example images during this stage to ensure it has been exposed to all possible scenarios. See Algorithm 1.
Multiple Image Extraction: For features in $\mathcal{F}_{m}$, the VLM generates code $\mathcal{G}_{m}$, which extends $\mathcal{G}_{s}$ to process sequential image-segmentation pairs $\{(I_{t-i},S_{t-i}),\dots,(I_{t},S_{t})\}$. One example of such a pair is provided to the VLM during generation. See Algorithm 2.
The final feature extraction code $\mathcal{G}$ is defined in Eq. 2:
$\mathcal{G}=\begin{cases}\mathcal{G}_{s},&\text{if }\mathcal{F}_{m}=\emptyset;\\ \mathcal{G}_{m}(\mathcal{G}_{s}),&\text{if }\mathcal{F}_{m}\neq\emptyset.\end{cases}$ (2)
Algorithm 1 Generating Extraction Code: Single Image
1:Input: Feature set $\mathcal{F}$, images $\{\mathcal{I}_{i}\}_{i=1}^{n}$, VLM $\mathcal{M}$
2:Initialize $\mathcal{G}_{s}=null$
3:$examples\leftarrow\emptyset$
4:for each image $\mathcal{I}_{i}$ in $\{\mathcal{I}_{i}\}_{i=1}^{n}$ do
5:   $\mathcal{S}_{i}\leftarrow\text{Seg}(\mathcal{I}_{i})$
6:   Add $(\mathcal{I}_{i},\mathcal{S}_{i})$ to $examples$
7:end for
8:$\mathcal{G}_{s}\leftarrow\mathcal{M}_{\text{Generate}}(\mathcal{F},examples)$
Algorithm 2 Generating Extraction Code: Image List
1:Input: Feature set $\mathcal{F}$, images $\{\mathcal{I}_{i}\}_{i=1}^{n}$, $\mathcal{G}_{s}$, VLM $\mathcal{M}$
2:Initialize $\mathcal{G}_{m}=null$
3:$examples\leftarrow\emptyset$
4:for each image $\mathcal{I}_{i}$ in $\{\mathcal{I}_{i}\}_{i=1}^{n}$ do
5:   $\mathcal{S}_{i}\leftarrow\text{Seg}(\mathcal{I}_{i})$
6:   Add $(\mathcal{I}_{i},\mathcal{S}_{i})$ to $examples$
7:end for
8:$\mathcal{F}_{m}\leftarrow{\mathcal{M}_{\text{JudgeMultiFrame}}}(\mathcal{F})$
9:if $\mathcal{F}_{m}\neq\emptyset$ then
10:   $\mathcal{G}_{m}\leftarrow\mathcal{M}_{\text{Generate}}(\mathcal{F},examples,\mathcal{G}_{s})$
11:end if
12:Return: $\mathcal{G}_{m}$
Step 3: Dataset Generation and Training an Extractor CNN. In this step, we execute $\mathcal{G}$ on trajectories collected from the environment to generate the feature dataset. For each trajectory, a series of images $\mathcal{T}=\{I_{1},I_{2},\dots,I_{k}\}$ is generated, with corresponding visual segmentations $\mathcal{S}_{t}=\text{Seg}(I_{t})$ for each $t$. The extractor code $\mathcal{G}$ processes each image-segmentation pair $(I_{t},S_{t})$ independently to extract features $\mathbf{v}_{t}=\{v_{1},v_{2},\dots,v_{p}\}$, where $v_{i}\in\mathbb{R}$.
The dataset, consisting of image-feature pairs $\{(\mathcal{I}_{t},\mathbf{v}_{t})\}$, is then used to train an Extractor CNN as a batch-operable and efficient replacement to $\mathcal{G}_{s}$. The trained model predicts semantic features directly from raw images, enabling significantly faster feature extraction during RL training.
If the features require temporal context (i.e., $\mathcal{F}_{m}\neq\emptyset$), $\mathcal{G}_{m}$, consisting of simple mathematical operations on the outputs of $\mathcal{G}_{s}$, is manually adapted to utilize the Extractor CNN in place of $\mathcal{G}_{s}$. For example, in the Skiing environment, $\mathcal{G}_{m}$ retrieves the static features $\mathcal{F}_{s}$ by directly obtaining the values from $\mathcal{G}_{s}$ applied to the last image and acquires the only temporal feature: skiier horizontal speed through subtracting the skiier positions between the first and last image frames. The manual adaptation preserves the exact same logic while ensuring that the subtraction operation is properly batched, significantly accelerating the training process by several orders of magnitude. Without this adaptation, training would be prohibitively time-consuming.
Discussion. Recognizing that hallucination is a well-documented challenge in generative VLM models [33], we provide segmentation along with raw images in Step 2 and 3, using a pre-defined segmentation function $Seg$ that detects objects and their positions in an image by contour and color analysis (details provided in Appendix VII-B). This strategy aligns with common practices in the field [34, 35]. By incorporating segmentation, we mitigate the risk of hallucination and ensure that the generated features and codes are grounded in accurate visual data.
TABLE I: Prompting statistics for VLM-generated code. “Code failure” indicates code that cannot import or run. “Excluded as constant” refers to code that generates invariant outputs across diverse samples and is therefore excluded. “Majority selection coverage” represents the proportion of code yielding identical outputs in the majority voting process. Step 1 refers to the process of generating $\mathcal{G}_{s}$, while Step 2 refers to the process of generating $\mathcal{G}_{m}$.
Environment Type Code Failure Rate (%) Excluded as Constant (%) Majority Selection Coverage (%)
Cart-pole Step 1 0 0 38
Cart-pole Step 2 0 0 100
Skiing Step 1 28 14 26
Skiing Step 2 0 0 100
Boxing Step 1 20 46 14
Boxing Step 2 - - -
TABLE II: Performance of Extractor CNN in extracting feature vectors across different environments, compared with ground truth from VLM-generated code. Continuous components are evaluated with MSE, normalized to [0,1], and scaled by $10^{-3}$; boolean components are evaluated with accuracy. Dashes indicate vacant entries due to varying feature lengths and types across environments.
Environment Feature Vector Indices (1–11)
1 2 3 4 5 6 7 8 9 10 11
Skiing (MSE) 0.11 0.02 0.16 0.20 1.06 2.13 1.93 0.50 1.81
Cart-pole (MSE) 0.008 0.016
Boxing (MSE) 1.07 2.30 23.25 19.42 6.80 0.98 21.71 20.26 12.12
Boxing (Accuracy) 0.993 0.990
III-B Tree-based RL Training
After obtaining the Extractor CNN, we train an interpretable tree-based model with RL to connect the semantic features with action outputs. During RL training, our framework utilizes the pre-trained Extractor CNN to extract semantic features from image observations, as shown in Fig. 1. The policy network takes these semantic features as input and predicts a distribution over actions, from which an action is sampled and applied to the controlled agent at each time step. Unlike traditional RL algorithms that use black-box models as the policy network, we use an Interpretable Control Tree (ICT) as our policy network to achieve transparent decision-making and human-aligned reasoning while remaining differentiable for end-to-end optimization.
Interpretable Control Tree: Interpretable Control Tree (ICT) is adapted from Interpretable Continuous Control Tree (ICCT) proposed by Paleja et al.  [6]. The ICT is structured similarly as ICCT, where each decision node determines the routing of input features, $\mathbf{x}$, to sparse linear controllers at the leaf nodes.
Each decision node, $i$, is parameterized by weights $\mathbf{w}_{i}$, a bias $b_{i}$, and a steepness parameter $\alpha$ to output a fuzzy Boolean in $[0,1]$, as shown in Eq. 3.
$y_{i}=\sigma\left(\alpha\left(w_{i}^{k}x^{k}-b_{i}\right)\right)$ (3)
The most important feature, $k$, of node, $i$, is selected in Eq. 4 by identifying the weight with the largest magnitude
$k=\arg\max_{j}\left|w_{i}^{j}\right|$ (4)
Since the $\arg\max$ operator is non-differentiable, we approximate this function with the Gumbel-Softmax  [36], as per Eq. 5, where $g_{j}\sim\text{Gumbel}(0,1)$ and $\tau$ is a temperature parameter.
$f(w_{i})^{k}=\frac{\exp\left(\frac{w_{i}^{k}+g_{k}}{\tau}\right)}{\sum_{j}\exp\left(\frac{w_{i}^{j}+g_{j}}{\tau}\right)}$ (5)
During inference, a hard threshold is applied to enforce discrete decision-making using only one feature per node, as per Eq. 6.
$y_{i}=\mathbbm{1}(\alpha(w_{i}^{k}x^{k}-b_{i})>0)$ (6)
While during training, gradients are allowed to flow through discrete decisions using the straight-through trick [37]. This approach ensures that each decision node remains interpretable while allowing for end-to-end gradient-based learning.
Each leaf node, $\ell_{d}$, in the ICT is parameterized by a set of sparse linear controllers, as formulated in Eq. 7, where $\mathbf{\beta}_{d}$ is a sparsely activated weight vector, $\phi_{d}$ is a bias term, and $\mathbf{u}_{d}$ is a $k$-hot encoding vector that enforces sparsity, computed in Eq. 8. The per-leaf selector weight, $\theta_{d}$, learns the relative importance of different input features. The function $h(\cdot)$ provides a differentiable approximation of the $k$-hot function by assigning a value of 1 to the top $k$ elements in $|\mathbf{\theta}_{d}|$ (based on magnitude) and 0 to the rest. Differentiability is preserved using the straight-through trick [37].
$PMF/PDF=(\mathbf{u}_{d}\circ\mathbf{\beta}_{d})^{\top}(\mathbf{u}_{d}\circ\mathbf{x})+\mathbf{u}_{d}^{\top}\phi_{d}$ (7)
$\mathbf{u}_{d}=h\left(\mathrm{softmax}\left(|\mathbf{\theta}_{d}|\right)\right),$ (8)
Unlike ICCT, ICT naturally handles continuous and discrete action space. In the discrete case, leaf nodes in ICT output the possibility of taking each action as a Probability Mass Function (PMF). If there are six possible actions in the discrete action space, then there will be six sparse linear functions of the input features followed by a softmax function at each leaf node. And for the continuous action space, leaf nodes output the Probability Density Function (PDF) over possible actions.
IV Experiments
IV-A Environment
We evaluate our approach on two Atari environments: Skiing and Boxing, and a classic control domain: Cart-pole, where we use image snapshots instead of the built-in feature vector as the observation space. The OpenAI Atari platform is preferable because all their environments provide raw image state spaces without any pre-defined feature vectors or feature sets. This ensures that the VLM engages in genuine reasoning over the environment to identify key semantic features, rather than relying on prior internet-based knowledge, a phenomenon referred to as training data memorization [38].
IV-B Feature Extraction
IV-B1 VLM Deducing Key Semantic Features.
For this step, we utilize GPT-4o, the current state-of-the-art in language reasoning. Across all three environments, the VLM consistently produces a deterministic set of semantic features for decision-making, achieving performance comparable to human reasoning. For all three environments, we query the VLM 50 times and use majority vote to get the most prominent answer from the VLM. We present the identified feature set for the Skiing environment in Fig. 1. Other domains are included in Appendix VII-D.
(a) Boxing
(b) Cartpole
(c) Skiing
Figure 2: Results across three domains. Higher scores are better.
IV-B2 VLMs for Feature Extraction Code
For this step, we utilize claude-3.5-sonnet-20241022, the current state-of-the-art in code generation. To address common challenges with VLM-generated code, such as non-functionality or runtime errors, we query the VLM 50 times using the same prompt (temperature = 0 for reproducibility) and select the most reliable feature extraction code through majority voting.
The generated codes are evaluated on a randomly collected dataset of 100 images and filtered based on the following criteria: (1) the code must compile and execute successfully and (2) it must produce diverse feature outputs rather than constant or repetitive values across the dataset. Candidates producing identical outputs are grouped into clusters, and the largest cluster is chosen as the final feature extraction code. Table I presents statistics from this process, including code failure rates, exclusions due to constant outputs, and majority selection coverage. The majority selection coverage appears low because features related to relative distances are normalized differently, while features related to object positions remain identical.
It is important to note that while Step 1, which generates $\mathcal{G}_{s}$ for feature extraction from a single image frame, is essential for all tasks, Step 2, which generates $\mathcal{G}_{m}$ for feature extraction across multiple image frames, is only necessary when the task’s features depend on temporal information. In the Boxing environment, the feature space does not depend on temporal inference, making Step 2 unnecessary. Consequently, Step 2 is marked as empty for Boxing in Table I. However, the VLM was still queried 50 times, and in all 50 cases, it successfully determined that no code generation was needed, achieving a 100% success rate under this metric.
IV-B3 Extractor CNN Performance Analysis
In Step 3 of our methodology, we address the inefficiencies of the VLM-generated feature extraction code for single-image frames, $\mathcal{G}_{s}$, by training an Extractor CNN as its replacement. While $\mathcal{G}_{s}$ can be executed directly on the dataset, the processing pipeline is inherently non-batchable, as it handles image segmentations independently (see Appendix VII-B for examples). This non-batchability severely impacts RL training, causing slowdowns of 200–2000× compared to batchable operations (Extractor CNN), as demonstrated in Table III. From practical experience, this would translate to training times exceeding multiple weeks on an H40-level GPU for a single policy to converge, making this infeasible.
To constrain computing time to a reasonable range, we train an Extractor CNN to replace $\mathcal{G}_{s}$ and efficiently compute $\mathcal{F}_{s}$, the features derived from a single image frame. The Extractor CNN is specifically designed to resolve the slowdown caused by segmentation-based analysis in $\mathcal{G}_{s}$. For each environment, we train an Extractor CNN on 7,000 images collected from diverse trajectories, validate it on 2,000 images, and test it on 1,000 images. The models achieved a mean squared error (MSE) consistently below 0.02 for each feature in the output vector when compared to the VLM-generated ground truth. Table II provides detailed performance statistics. These results confirm that the Extractor CNN is a reliable and computationally efficient replacement for $\mathcal{G}_{s}$.
TABLE III: Runtime speedup for processing a batch of images across different environments on the same device (averaged over 10 iterations). Results denoted with * are derived from a different device with a better gpu.
Env. Batch VLM (s) CNN (s) Speedup
Skiing 8 0.870 0.011 81.02x
16 1.745 0.003 523.37x
32 3.523 0.005 743.77x
64 7.055 0.008 877.90x
128 14.449 0.021 693.74x
Cart-pole 8 12.726 0.017 738.57x
16 25.679 0.040 637.06x
32 50.204 0.066 763.96x
64 99.660 0.111 898.09x
128* 254.484 0.097 2617.22x
Boxing 8 0.400 0.017 23.07x
16 0.785 0.004 191.03x
32 1.581 0.007 238.51x
64 3.146 0.009 344.34x
128 6.240 0.025 254.40x
IV-C Training Results
Figure 3: Visualization of ICT in the Skiing Environment.
Figure 4: Demonstration of the discrepancy between the physics simulation and the visualization.
We train ICT, our proposed interpretable and verifiable method that learns a tree structure with RL, alongside several baseline models across all environments. Below is an explanation of each baseline model used in the evaluation:
  • CNN: An uninterpretable black-box model that maps raw visual inputs directly to the action space, trained with RL.
  • VLM-ICT: Our proposed interpretable and verifiable method, using an ICT to connect the semantic feature space with the action space, trained with RL.
  • VLM-ICT (Rand Init): Our proposed method, but initialized randomly and not trained with RL.
  • VLM-DT Heuristic: A baseline where VLM generates a heuristic tree to connect the semantic feature space with the action space.
  • VLM-MLP: An uninterpretable benchmark that connects a multi-layer perceptron between the semantic feature space and the action space.
  • Random Policy: A policy that selects actions by randomly sampling from the action space.
Across all three environments, VLM-ICT consistently outperforms other uninterpretable models (MLP) built on the semantic feature space, and demonstrates performance close to the CNN+RL black-box benchmark representing the upper limit of task performance. Notably, in the Cart-pole environment, VLM-ICT even outperforms the black-box CNN model. This is because the CNN fails to accurately capture the pole’s angle, a critical feature for formulating an effective policy. These consistent results highlight the robustness and effectiveness of our proposed method.
The learned ICT in Skiing domain is shown in Fig. 3. This tree was generated fully autonomously using the VLM $\rightarrow$ CNN $\rightarrow$ RL pipeline. The learned trees (ICT) are interpretable both subjectively and objectively, provided they remain small enough, as demonstrated in a previous user study [6]. For the visualization of our learned ICT in Boxing and Cart-pole domains, please refer to Appendix VII-A.
V Limitations and Future Works
In the Cart-pole environment, we observe that the performance of trained VLM-ICT and VLM-MLP has yet to match that of RL policies trained using ground truth features [39, 6]. This discrepancy stems from the gap between the underlying physics simulation and its visual representation, as illustrated in Fig. 4. While the trained Extractor CNN’s predictions closely approximate the feature vectors extracted by the VLM-generated code, they still exhibit deviations from the ground truth feature vectors. This highlights a fundamental challenge in semantic RL from visual observations, where the extracted features may not perfectly align with the underlying physical state. Despite this limitation, our framework still outperforms the CNN-based policy, demonstrating the promise of VLMs in RL. To the best of our knowledge, we are the first to attempt solving the Cart-pole problem using image-based observations, demonstrating a novel and promising direction for interpretable RL.
Future works could explore RL-based backpropagation to fine-tune the VLM models, enabling large language models to iteratively refine their outputs by learning from feedback generated through interactions with RL agents [40]. This approach has the potential to enable VLMs to identify discrepancies between the underlying physics simulation and its visual representation, leading to improved feature extraction for downstream RL tasks.
VI Conclusion
We introduce SILVA, a novel framework for achieving automated semantic interpretability in RL. It automates feature space construction and extraction with zero-shot VLM querying, making it generalizable to unseen environments without human-in-the-loop. To overcome VLMs’ inability to generate valid decision policies, SILVA leverages an ICT for RL training, obtaining transparent, human-understandable, and verifiable decision rules. The framework outperforms black-box models built on the same feature space and achieves performance on par with purely uninterpretable black-box models. While all modern interpretable models, including tree-based approaches, face challenges in complex environments, SILVA stands out among them as the first to achieve automated semantic interpretability in RL, marking a significant advancement that underscores its impact.
Ethics and Impact Statement
SILVA enhances interpretability in RL by automating semantic feature extraction via VLMs, reducing reliance on human annotations and fostering transparent AI decision-making. By structuring policies in an interpretable tree-based format, SILVA aligns with ethical AI principles emphasizing explainability, accountability, and trust.
However, reliance on VLMs introduces biases from pre-trained models, potentially causing discriminatory behavior or unreliable feature extraction. While SILVA improves interpretability, it still depends on black-box models for feature generation, challenging full transparency.
To mitigate these risks, future work could incorporate bias audits, human-in-the-loop verification, and standardized fairness metrics.

### v2

@@
 Abstract
-Semantic Interpretability in Reinforcement Learning (RL) enables transparency, accountability, and safer deployment by making the agent’s decisions understandable and verifiable.
-Achieving this, however, requires a feature space composed of human-understandable concepts, which traditionally rely on human specification and fail to generalize to unseen environments.
-In this work, we introduce Semantically Interpretable Reinforcement Learning with Vision-Language Models Empowered Automation (SILVA), an automated framework that leverages pre-trained vision-language models (VLM) for semantic feature extraction and interpretable tree-based models for policy optimization.
-SILVA first queries a VLM to identify relevant semantic features for an unseen environment, then extracts these features from the environment.
-Finally, it trains an Interpretable Control Tree via RL, mapping the extracted features to actions in a transparent and interpretable manner.
-To address the computational inefficiency of extracting features directly with VLMs, we develop a feature extraction pipeline that generates a dataset for training a lightweight convolutional network, which is subsequently used during RL.
-By leveraging VLMs to automate tree-based RL, SILVA removes the reliance on human annotation previously required by interpretable models while also overcoming the inability of VLMs alone to generate valid robot policies, enabling semantically interpretable reinforcement learning without human-in-the-loop.
-I Introduction
+Semantic interpretability in Reinforcement Learning (RL) enables transparency and verifiability by making the agent’s decisions understandable and verifiable.
+Achieving this, however, requires a feature space composed of human-understandable concepts, which traditionally rely on human specification and may fail to generalize to unseen environments.
+We introduce interpretable Tree-based Reinforcement learning via Automated Concept Extraction (iTRACE), an automated framework that leverages pre-trained vision-language models (VLM) for semantic feature extraction and interpretable tree-based models for policy optimization.
+iTRACE first extracts semantically meaningful features, then maps them to policies via interpretable trees.
+To address the impracticality of running VLMs in RL loops, we distill their outputs into a lightweight model.
+By leveraging Vision-Language Models (VLMs) to automate tree-based reinforcement learning, iTRACE eliminates the need for human annotation traditionally required by interpretable models, while also addressing the limitations of VLMs alone, such as their lack of grounding in action spaces and inability to directly optimize policies.
+iTRACE outperforms MLP baselines that use the same interpretable features and matches the performance of CNN-based policies, producing verifiable, semantically interpretable, and human-aligned behaviors without requiring human annotation.
+1 Introduction
 Reinforcement learning (RL) agents operate in dynamic environments, often developing policies that rely on latent representations not directly interpretable by humans [28].
@@
 Decision trees have been used to represent policies in a structured, human-readable format [30, 26, 18].
-Similarly, rule-based methods, such as logical programs and symbolic reasoning approaches, offer an alternative by encoding policies in a form that can be explicitly followed and verified [7, 8].
-Additionally, hierarchical decision-making frameworks have been explored, where high-level policies generate interpretable subgoals while low-level controllers handle execution [9, 10].
+Rule-based methods offer an alternative by encoding policies in a form that can be explicitly followed and verified [7, 12].
 However, none of these prior works address high-dimensional image or video inputs; instead, they rely either on human-defined semantic concepts or ground-truth feature vectors provided by the environment.
@@
 These models, known for their capabilities in fine-grained visual understanding, language reasoning, and decision-making  [4, 38, 6], have seen widespread use in various tasks  [38, 39].
-In this work, we demonstrate that VLMs can substitute humans in defining and extracting semantic features, serving not merely as a “scarecrow in oz”  [22] or a cost-effective fallback to human labor, but as a scalable and automated solution that matches human performance.
-We propose Semantically Interpretable Reinforcement Learning with Vision-Language Models Empowered Automation (SILVA), which is to the best of our knowledge the first automated framework that achieves semantic interpretability in RL and generalizes effectively to unseen environments.
+In this work, we demonstrate that VLMs can substitute humans in defining and extracting semantic features, serving not merely as a cost-effective fallback to human labor, but as a scalable and automated solution that matches human performance.
+We propose interpretable Tree-based Reinforcement learning via Automated Concept Extraction (iTRACE), which is to the best of our knowledge the first automated framework that achieves semantic interpretability in RL and generalizes effectively to unseen environments.
 By leveraging VLMs to extract features from raw environment inputs and training with Interpretable Control Trees, we manage to provide the rationale behind a RL algorithm in terms of human-understandable concepts.
-Notably, our framework (i) maintains competitive performance with black-box models while surpassing uninterpretable models built on the same feature space, (ii) supports both image and video inputs, effectively capturing temporal information, (iii) generalizes to unseen environments without human-in-the-loop, and (iv) enables computational efficiency by training a lightweight convolutional network as an alternative to VLM-based feature extraction.
-II Related Work
-II-A Interpretable Reinforcement Learning
-Figure 1: Overview of SILVA, our proposed framework.
+Notably, our framework (i) maintains competitive performance with black-box models while surpassing uninterpretable models built on the same feature space, (ii) supports both image and video inputs, effectively capturing temporal information, (iii) generalizes to unseen environments without human-in-the-loop, and (iv) enables computational efficiency by training a lightweight vision model as an alternative to VLM-based feature extraction.
+2 Related work
+2.1 Interpretable reinforcement learning
+Figure 1: Overview of iTRACE, our proposed framework.
 A.
 Feature Extraction involves a three-step process where: (1) The VLM identifies key features given domain knowledge and environment observation.
-(2) The VLM generates code to extract features from a single image frame and assesses whether certain key features require multiple frames.
-If so, it builds upon the single-frame code to extract these features.
-(3) Input-output pairs are collected using the extractor code and are used to train an Extractor Convolutional Neural Network (Extractor CNN) as a computationally efficient substitute.
+(2) The VLM generates code to extract features from a segmentation of a single image frame and determines whether certain key features require multiple frames.
+If so, it extends the single-frame extraction code to handle multi-frame inputs.
+The segmentation is performed using the Segment Anything Model (SAM).
+(3) An exploration policy is used to generate a dataset of feature–raw image frame pairs by applying SAM and the feature extraction code.
+To avoid the computational overhead of running SAM at every reinforcement learning step, we train an Extractor Module as an efficient substitute.
 B.
-Tree-based RL Training leverages the Extractor CNN to obtain semantic features from image observations, which are subsequently used as inputs to the policy network (an Interpretable Control Tree) to generate control actions for the RL agent.
-Interpretable RL is inherently challenging compared to standard supervised learning, as it involves long-term consequences to actions and latent receptions of data through exploration.
-[2] A significant line of research has explored interpretability in RL through generating visual saliency maps using Jacobian-based, perturbation-based, and attribution-based methods [23, 13, 24].
-However, these approaches primarily focus on post-hoc explanations and fall out of a stricter definition of Interpretable RL  [3], which requires all components including the input space and the decision-making process to be intelligible.
-Approaches to achieve intelligible decision-making includes interpretable tree-structures, rule and decision lists, and other rule-based methods [3], among which decision trees are arguably the easiest to verify  [25].
-Among state-of-the-art tree methods, Differentiable Decision Trees (DDTs) employ a Decision Tree structure that allows for online learning via stochastic gradient descent [25], while Interpretable Continuous Control Trees (ICCTs) extend this capability to continuous control tasks, ensuring both high performance and interpretability [26].
+Tree-based RL Training leverages the pretrained Extractor Module to obtain interpretable features from image observations, which are subsequently used as inputs to the policy network (an Interpretable Control Tree) to generate control actions for the RL agent.
+Interpretable RL is inherently challenging compared to standard supervised learning, as it needs to account for the long-term consequences of actions, partial observability, and sparse feedback during exploration [22].
+Post-hoc methods like Jacobian-based, perturbation-based, and attribution-based saliency maps  [25, 10, 34] offer limited interpretability, as they lack transparency in both inputs and decision-making [9].
+Approaches to achieve intelligible decision-making include rule lists and decision trees [9], among which decision trees are arguably the easiest to verify  [27].
+Among state-of-the-art tree methods, Differentiable Decision Trees (DDTs) support gradient-based optimization of interpretable models [27], while Interpretable Continuous Control Trees (ICCTs) extend this capability to continuous control tasks, balancing performance with interpretability [19].
 Similarly, POETREE (Policy Extraction through Decision Trees) [17] introduces dynamically growing probabilistic decision trees that adapt throughout training, allowing the tree complexity to evolve based on the task requirements.
-Our work addresses the limitation of prior tree-based methods by automating feature definition and extraction, thus enabling generalization to unseen environments.
-II-B Feature Extraction for Reinforcement Learning
-Common techniques that map high-dimensional inputs into low-dimensional feature spaces for RL includes linear nonparametric methods, such as manifold learning and spectral learning, and nonlinear parametric methods, such as deep learning  [12].
-Other work focuses on the extraction of visual features that are not inherently connected with semantic meanings.
-[23] employs a self-supervised interpretable network (SSINet) to produce fine-grained attention masks for highlighting task-relevant information.
-[13] computed perturbation-based saliency of the image input to deduce regions that had the most effect.
-However, only a limited number of studies have proposed approaches without human-in-the-loop to construct semantic feature spaces for RL training.
-[28] leverages information from Atari instruction manuals to construct the feature space, then uses unsupervised object detection and grounding to acquire the corresponding
-objects in the environment.
-However, they observe that current techniques exhibit a lack of reliability  [28].
-Our framework distinguishes from prior work by enabling automatic feature definition and extraction without human-in-the-loop.
-II-C VLMs for Reinforcement Learning
-VLMs have been applied in various aspects of RL training, including reward design, preference feedback, and representation deriving [29, 30, 20].
+Our work addresses the limitation of prior tree-based methods, which previously required manual feature engineering or were limited to structured inputs
+2.2 Feature extraction for reinforcement learning
+Nonparametric methods (e.g., manifold and spectral learning) and nonlinear parametric methods (e.g., deep learning) are commonly used in RL for feature extraction but typically do not produce interpretable features [15].
+Other work focuses on the extraction of visual features that improve performance but do not guarantee human-interpretable concepts.
+Yet, few works attempt to construct semantically meaningful feature spaces without human supervision for RL training.
+[35] leverages information from Atari instruction manuals to construct the feature space, then uses unsupervised object detection and grounding to acquire the corresponding objects in the environment.
+However, they report that the extracted features could be noisy and inconsistent  [35].
+Our framework differs from prior work by enabling automatic feature definition and extraction without human-in-the-loop.
+2.3 VLMs for reinforcement learning
+VLMs have been applied in various aspects of RL training, including reward design, preference feedback, and representation extraction [21, 33, 6].
 One prominent approach involves using VLMs to generate reward functions.
@@
 [6] extract promptable representations from VLMs and decode output tokens into embeddings, which are used to initialize RL policies.
-However, those representations are not derived for interpretability purposes and might not often be correct.
-To the best of our knowledge, our work is the first to utilize VLM for creating a semantic feature space for RL training.
-III Methodology
-We introduce Semantically Interpretable Reinforcement Learning with Vision-Language Models Empowered Automation (SILVA), a framework designed to automatically learn an interpretable decision-making RL policy from an unseen environment.
-SILVA leverages the visual comprehension and reasoning capabilities of VLMs automate feature extraction through a three-step pipeline: (1) it employs zero-shot prompting on the environment to identify a set of semantic features that deterministically contribute to the decision-making, (2) it generates feature extraction codes that captures both static and temporal information, and (3) it collects a dataset with VLM-generated extraction codes to train a lightweight convolutional neural network (Extractor CNN) as a computationally efficient substitute.
-Subsequently, utilizing the Extractor CNN, SILVA performs tree-based RL training with an Interpretable Control Tree, establishing a transparent mapping between the semantic feature space and action outputs to derive a semantically interpretable policy for decision-making.
-III-A Feature Extraction
-The feature extraction process of SILVA involves a three-step pipeline, as shown in Fig.
+However, these representations are not guaranteed to be accurate or semantically meaningful.
+To our knowledge, this is the first approach that uses VLMs to automatically construct semantically interpretable feature spaces for RL.
+3 Methodology
+We introduce interpretable Tree-based Reinforcement learning via Automated Concept Extraction (iTRACE), a framework designed to automatically learn an interpretable decision-making RL policy from an unseen environment.
+iTRACE leverages the visual comprehension and reasoning capabilities of VLMs automate feature extraction through a three-step pipeline: (1) it employs zero-shot prompting on the environment to identify a set of semantic features that deterministically contribute to the decision-making, (2) it generates feature extraction codes that captures both static and temporal information, and (3) it collects a dataset with VLM-generated extraction codes to train a lightweight vision model (which we refer to as the Extractor Module) as a computationally efficient substitute.
+Subsequently, utilizing the Extractor Module to replicate VLM feature extraction, iTRACE performs tree-based RL training with an Interpretable Control Tree, establishing a transparent mapping between the semantic feature space and action outputs to derive a semantically interpretable policy for decision-making.
+3.1 Feature extraction
+The feature extraction process of iTRACE involves a three-step pipeline, as shown in Fig.
 1.
 In Step 1, we query the VLM to generate key features that contribute to the decision in the environment.
-In Step 2, we query the VLM to generate code that extracts the identified features from the environment, either from a single image or a list of images.
-In Step 3, we collect a dataset from the environment with the VLM-generated extractor code, with which we train an Extractor CNN that will be used in RL training to enable batch computing and increase program speed.
+In Step 2, we query the VLM to generate code that extracts the identified features, either from a single image or a list of images depending on whether the features demand temporal context.
+Both raw observations and segmentation masks are provided to the VLM in this process.
+In Step 3, we collect a dataset from the environment with the VLM-generated extractor code, with which we train an Extractor Module to increase program speed in RL training.
 The methodology is formalized below.
@@
   • $\mathcal{D}$: Basic domain knowledge about the environment.
-  • $\mathcal{I}_{0}$: An example image of the environment.
+  • $\mathcal{I}_{0}$: A diverse list of example image observations from the environment.
 The VLM identifies key semantic features critical for decision-making, represented as two distinct sets:
@@
 $\mathcal{F}=\mathcal{F}_{s}\cup\mathcal{F}_{m},\quad\mathcal{F}=\mathcal{M}(\mathcal{D},\mathcal{I}_{0}).$ (1)
-For example, in the Skiing environment, $\mathcal{F}_{s}$ includes features such as the position of skier, the position of next gate, the distance to next gate, the position of nearest tree, and the distance to nearest tree, while $\mathcal{F}_{m}$ includes the current horizontal speed of the skiier.
-This example is also included in Fig.
-1.
+For example, in the Skiing environment, $\mathcal{F}_{s}$ includes the positions of the next gate (left pole x, right pole x, and y), the closest tree (x, y) and the skier’s x-position, while $\mathcal{F}_{m}$ includes the horizontal speed of the skier.
+We instruct the VLM to provide a "requires_temporal_context" field for each feature, enabling it to automatically distinguish between static and temporal features and assign them to $\mathcal{F}_{s}$ or $\mathcal{F}_{m}$.
+While $\mathcal{F}_{m}$ may be unnecessary in some environments, it is essential in others.
+Returning to the Skiing example, the skier must actively tilt left and right to maintain speed.
+Without knowledge of the skier’s motion direction, it is ambiguous which way the skier should tilt.
+We allow the VLM to determine whether temporal context is required for effective decision-making in a given environment.
+A complete list of extracted features is provided in Appendix A.2.2.
 Step 2: Querying the VLM for Feature Extraction Code.
-In this step, we query the VLM to generate feature extraction code, denoted by $\mathcal{G}$, to retrieve the identified set of features $\mathcal{F}$.
-To mitigate hallucination, we provide visual segmentations for each image, which will be covered in the discussion section.
-The visual segmentation $\mathcal{S}$ is computed using a segmentation function, represented as $Seg$, such that $\mathcal{S}=\text{Seg}(\mathcal{I})$.
-The code generation process involves two stages:
-Single Image Extraction: For features in $\mathcal{F}$, the VLM generates code $\mathcal{G}_{s}$ to extract as many features as possible from a single image-segmentation pair $(\mathcal{I},\mathcal{S})$.
-In practice, as demonstrated in our experiments, this approach effectively recovers all features in $\mathcal{F}_{s}$.
-To enhance the robustness of the generated code, the VLM is provided with multiple example images during this stage to ensure it has been exposed to all possible scenarios.
+In this step, we query the VLM to generate feature extraction code, denoted by $\mathcal{G}$, to retrieve the identified set of features $\mathcal{F}$ from one or multiple image-segmentation pairs $(\mathcal{I},\mathcal{S})$.
+Recognizing that hallucination is a well-documented challenge in generative VLM models [16] (see Appendix.
+A.5), we use segmentation to provide accurate spatial information and ground the generated extraction code in accurate visual data.
+This strategy aligns with common practices in the field [20, 37].
+The visual segmentation $\mathcal{S}$ is computed using Segment Anything [13], denoted by $Seg$, such that $\mathcal{S}=\text{Seg}(\mathcal{I})$.
+We use the ViT-H SAM model with its default parameters and no modifications to ensure the generalizability of our approach.
+Example Segmentations Labeling: While segmentation improves spatial grounding, its interpretation by the VLM remains prone to hallucination.
+To ensure the generated code generalizes across diverse environment frames, the VLM must be exposed to a representative set of example images.
+However, the raw JSON format of segmentation can lead to confusion between segments across different images.
+To address this, we first prompt the VLM to label a small set of segmentations associated with example frames, processing one image at a time.
+An example of this labeling process is shown in Fig.
+9(b), where all segment labels are generated by the VLM based on raw segmentation masks.
+Following this, depending on whether the feature set requires temporal context, the VLM is instructed to take one or multiple image-segmentation pairs $(\mathcal{I},\mathcal{S})$.
+The labeled segmentations serve as reference examples, helping the VLM infer the most likely attribute values of relevant objects and formulate logic to filter out noisy segments.
+The code generation process involves two steps:
+Single Image Extraction: For features in $\mathcal{F}_{s}$, the VLM generates code $\mathcal{G}_{s}$ to extract their values from a single image-segmentation pair $(\mathcal{I},\mathcal{S})$.
+To enhance the robustness of the generated code, the VLM is provided with multiple example images and their labeled segmentation during this stage to ensure it has been exposed to most possible scenarios.
 See Algorithm 1.
-Multiple Image Extraction: For features in $\mathcal{F}_{m}$, the VLM generates code $\mathcal{G}_{m}$, which extends $\mathcal{G}_{s}$ to process sequential image-segmentation pairs $\{(I_{t-i},S_{t-i}),\dots,(I_{t},S_{t})\}$.
-One example of such a pair is provided to the VLM during generation.
+Also see Fig.
+9(c) for an example output on segmentation masks in the Skiing environment.
+Multiple Image Extraction: For features in $\mathcal{F}_{m}$, if any, the VLM generates code $\mathcal{G}_{m}$, which extends $\mathcal{G}_{s}$ to process sequential image-segmentation pairs $\{(I_{t-i},S_{t-i}),\dots,(I_{t},S_{t})\}$.
+Multiple image-segmentation pairs $(\mathcal{I},\mathcal{S})$ are provided to the VLM during generation.
 See Algorithm 2.
@@
 8:  $\mathcal{G}_{s}\leftarrow\mathcal{M}_{\text{Generate}}(\mathcal{F},examples)$
-Algorithm 2 Generating Extraction Code: Image List
+Algorithm 2 Generating Extraction Code: Multiple Images
 1:  Input: Feature set $\mathcal{F}$, images $\{\mathcal{I}_{i}\}_{i=1}^{n}$, $\mathcal{G}_{s}$, VLM $\mathcal{M}$
@@
 12:  Return: $\mathcal{G}_{m}$
-Step 3: Dataset Generation and Training an Extractor CNN.
+(a) Original image observation.
+(b) VLM-labeled segmentation.
+(c) VLM-extracted features.
+Figure 2: An example of how raw image observations are processed in our pipeline.
+(a) Raw image observations are segmented using Segment Anything.
+(b) A VLM assigns semantic labels to segments in a small set of example images.
+(c) A VLM then generates code that extracts target features from segmentation masks, applicable to all images.
+Step (b) is only performed on the example set to guide code generation.
+Step 3: Dataset Generation and Training the Extractor Module.
 In this step, we execute $\mathcal{G}$ on trajectories collected from the environment to generate the feature dataset.
 For each trajectory, a series of images $\mathcal{T}=\{I_{1},I_{2},\dots,I_{k}\}$ is generated, with corresponding visual segmentations $\mathcal{S}_{t}=\text{Seg}(I_{t})$ for each $t$.
-The extractor code $\mathcal{G}$ processes each image-segmentation pair $(I_{t},S_{t})$ independently to extract features $\mathbf{v}_{t}=\{v_{1},v_{2},\dots,v_{p}\}$, where $v_{i}\in\mathbb{R}$.
-The dataset, consisting of image-feature pairs $\{(\mathcal{I}_{t},\mathbf{v}_{t})\}$, is then used to train an Extractor CNN as a batch-operable and efficient replacement to $\mathcal{G}_{s}$.
+The extractor code $\mathcal{G}$ processes each image-segmentation pair $(I_{t},S_{t})$ independently to extract features $\mathbf{f}_{t}=\{f_{1},f_{2},\dots,f_{n}\}$.
+The dataset, consisting of image-feature pairs $\{(\mathcal{I}_{t},\mathbf{f}_{t})\}$, is then used to train an Extractor Module as a batch-operable and efficient replacement to $\mathcal{G}_{s}$.
 The trained model predicts semantic features directly from raw images, enabling significantly faster feature extraction during RL training.
-If the features require temporal context (i.e., $\mathcal{F}_{m}\neq\emptyset$), $\mathcal{G}_{m}$, consisting of simple mathematical operations on the outputs of $\mathcal{G}_{s}$, is manually adapted to utilize the Extractor CNN in place of $\mathcal{G}_{s}$.
-For example, in the Skiing environment, $\mathcal{G}_{m}$ retrieves the static features $\mathcal{F}_{s}$ by directly obtaining the values from $\mathcal{G}_{s}$ applied to the last image and acquires the only temporal feature: skiier horizontal speed through subtracting the skiier positions between the first and last image frames.
-The manual adaptation preserves the exact same logic while ensuring that the subtraction operation is properly batched, significantly accelerating the training process by several orders of magnitude.
-Without this adaptation, training would be prohibitively time-consuming.
-Discussion.
-Recognizing that hallucination is a well-documented challenge in generative VLM models [33], we provide segmentation along with raw images in Step 2 and 3, using a pre-defined segmentation function $Seg$ that detects objects and their positions in an image by contour and color analysis (details provided in Appendix VII-B).
-This strategy aligns with common practices in the field [34, 35].
-By incorporating segmentation, we mitigate the risk of hallucination and ensure that the generated features and codes are grounded in accurate visual data.
-TABLE I: Prompting statistics for VLM-generated code.
-“Code failure” indicates code that cannot import or run.
-“Excluded as constant” refers to code that generates invariant outputs across diverse samples and is therefore excluded.
-“Majority selection coverage” represents the proportion of code yielding identical outputs in the majority voting process.
-Step 1 refers to the process of generating $\mathcal{G}_{s}$, while Step 2 refers to the process of generating $\mathcal{G}_{m}$.
-Environment Type Code Failure Rate (%) Excluded as Constant (%) Majority Selection Coverage (%)
-Cart-pole Step 1 0 0 38
-Cart-pole Step 2 0 0 100
-Skiing Step 1 28 14 26
-Skiing Step 2 0 0 100
-Boxing Step 1 20 46 14
-Boxing Step 2 - - -
-TABLE II: Performance of Extractor CNN in extracting feature vectors across different environments, compared with ground truth from VLM-generated code.
-Continuous components are evaluated with MSE, normalized to [0,1], and scaled by $10^{-3}$; boolean components are evaluated with accuracy.
-Dashes indicate vacant entries due to varying feature lengths and types across environments.
-Environment Feature Vector Indices (1–11)
-1 2 3 4 5 6 7 8 9 10 11
-Skiing (MSE) 0.11 0.02 0.16 0.20 1.06 2.13 1.93 0.50 1.81
-Cart-pole (MSE) 0.008 0.016
-Boxing (MSE) 1.07 2.30 23.25 19.42 6.80 0.98 21.71 20.26 12.12
-Boxing (Accuracy) 0.993 0.990
-III-B Tree-based RL Training
-After obtaining the Extractor CNN, we train an interpretable tree-based model with RL to connect the semantic features with action outputs.
-During RL training, our framework utilizes the pre-trained Extractor CNN to extract semantic features from image observations, as shown in Fig.
+In our experiments, we use a CNN-based Extractor Module (architecture included in Appendix A.3) due to its low computational overhead in simulated environments.
+However, the framework is compatible with any vision model, including fine-tuned ViT transformers or other task-specific architectures based on deployment needs.
+If some features require temporal context (i.e., $\mathcal{F}_{m}\neq\emptyset$), $\mathcal{G}_{m}$, consisting of simple mathematical operations on the outputs of $\mathcal{G}_{s}$, is manually adapted to utilize the Extractor Module in place of $\mathcal{G}_{s}$.
+For example, in the Skiing environment, $\mathcal{G}_{m}$ calculates skier horizontal speed by subtracting the positions returned by $\mathcal{G}_{s}$ The manual adaptation preserves the exact same logic while ensuring that the subtraction operation is properly batched, significantly accelerating the training process by several orders of magnitude.
+3.2 Tree-based RL training
+After obtaining the Extractor Module, we train an interpretable tree-based model with RL to connect the semantic features with action outputs.
+During RL training, our framework utilizes the pre-trained Extractor Module to extract semantic features from image observations, as shown in Fig.
 1.
@@
 Unlike traditional RL algorithms that use black-box models as the policy network, we use an Interpretable Control Tree (ICT) as our policy network to achieve transparent decision-making and human-aligned reasoning while remaining differentiable for end-to-end optimization.
-Interpretable Control Tree: Interpretable Control Tree (ICT) is adapted from Interpretable Continuous Control Tree (ICCT) proposed by Paleja et al.
+3.2.1 Interpretable control tree
+Interpretable Control Tree (ICT) is adapted from Interpretable Continuous Control Tree (ICCT) proposed by Paleja et al.
 [18].
 The ICT is structured similarly as ICCT, where each decision node determines the routing of input features, $\mathbf{x}$, to sparse linear controllers at the leaf nodes.
-Each decision node, $i$, is parameterized by weights $\mathbf{w}_{i}$, a bias $b_{i}$, and a steepness parameter $\alpha$ to output a fuzzy Boolean in $[0,1]$, as shown in Eq. 3.
+Each decision node, $i$, is parameterized by weights $\mathbf{w}_{i}$, a bias $b_{i}$, and a steepness parameter $\alpha$ to output a fuzzy Boolean in $[0,1]$, as defined in Eq. 3.
 $y_{i}=\sigma\left(\alpha\left(w_{i}^{k}x^{k}-b_{i}\right)\right)$ (3)
@@
 During inference, a hard threshold is applied to enforce discrete decision-making using only one feature per node, as per Eq. 6.
-$y_{i}=\mathbbm{1}(\alpha(w_{i}^{k}x^{k}-b_{i})>0)$ (6)
+$y_{i}=\mathbf{1}(\alpha(w_{i}^{k}x^{k}-b_{i})>0)$ (6)
 While during training, gradients are allowed to flow through discrete decisions using the straight-through trick [3].
@@
 $\mathbf{u}_{d}=h\left(\mathrm{softmax}\left(|\mathbf{\theta}_{d}|\right)\right),$ (8)
-Unlike ICCT, ICT naturally handles continuous and discrete action space.
+Unlike ICCT, ICT supports both discrete and continuous actions without re-architecture, which helps generalizability.
 In the discrete case, leaf nodes in ICT output the possibility of taking each action as a Probability Mass Function (PMF).
@@
 And for the continuous action space, leaf nodes output the Probability Density Function (PDF) over possible actions.
-IV Experiments
-IV-A Environment
-We evaluate our approach on two Atari environments: Skiing and Boxing, and a classic control domain: Cart-pole, where we use image snapshots instead of the built-in feature vector as the observation space.
-The OpenAI Atari platform is preferable because all their environments provide raw image state spaces without any pre-defined feature vectors or feature sets.
-This ensures that the VLM engages in genuine reasoning over the environment to identify key semantic features, rather than relying on prior internet-based knowledge, a phenomenon referred to as training data memorization [38].
-IV-B Feature Extraction
-IV-B1 VLM Deducing Key Semantic Features.
-For this step, we utilize GPT-4o, the current state-of-the-art in language reasoning.
+4 Experiments
+4.1 Environment
+We evaluate our approach across three diverse environments: Atari Skiing, a fast-paced pixel-art game requiring gate navigation; BabyAI-GoToRedBall, a symbolic grid world for instruction following; and OpenAI Gym Highway, a 2D environment for high-level autonomous driving decisions like lane switching and speed control.
+The Atari Skiing environment provides raw image state spaces without any pre-defined feature vectors or feature sets.
+BabyAI and Highway offer default feature vectors, but these vectors describe the full state of the environment rather than focusing on the task-relevant semantic features needed for decision making.
+By selecting these environments, we ensure that the VLM engages in genuine reasoning over the environment to identify key semantic features, rather than relying on prior internet-based knowledge, a phenomenon referred to as training data memorization [5].
+4.2 Feature extraction
+4.2.1 VLM deducing key semantic features.
+For this step, we utilize ChatGPT-4o-latest, one of the current leading models in language reasoning.
 Across all three environments, the VLM consistently produces a deterministic set of semantic features for decision-making, achieving performance comparable to human reasoning.
 For all three environments, we query the VLM 50 times and use majority vote to get the most prominent answer from the VLM.
-We present the identified feature set for the Skiing environment in Fig.
-1.
-Other domains are included in Appendix VII-D.
-(a) Boxing
-(b) Cartpole
-(c) Skiing
-Figure 2: Results across three domains. Higher scores are better.
-IV-B2 VLMs for Feature Extraction Code
-For this step, we utilize claude-3.5-sonnet-20241022, the current state-of-the-art in code generation.
+A complete set of those features are included in Appendix A.2.2.
+4.2.2 VLM generating Feature Extraction code.
+For this step, we utilize claude-3.5-sonnet-20241022, the current leading model in code generation.
 To address common challenges with VLM-generated code, such as non-functionality or runtime errors, we query the VLM 50 times using the same prompt (temperature = 0 for reproducibility) and select the most reliable feature extraction code through majority voting.
@@
 Candidates producing identical outputs are grouped into clusters, and the largest cluster is chosen as the final feature extraction code.
-Table I presents statistics from this process, including code failure rates, exclusions due to constant outputs, and majority selection coverage.
-The majority selection coverage appears low because features related to relative distances are normalized differently, while features related to object positions remain identical.
-It is important to note that while Step 1, which generates $\mathcal{G}_{s}$ for feature extraction from a single image frame, is essential for all tasks, Step 2, which generates $\mathcal{G}_{m}$ for feature extraction across multiple image frames, is only necessary when the task’s features depend on temporal information.
-In the Boxing environment, the feature space does not depend on temporal inference, making Step 2 unnecessary.
-Consequently, Step 2 is marked as empty for Boxing in Table I.
-However, the VLM was still queried 50 times, and in all 50 cases, it successfully determined that no code generation was needed, achieving a 100% success rate under this metric.
-IV-B3 Extractor CNN Performance Analysis
-In Step 3 of our methodology, we address the inefficiencies of the VLM-generated feature extraction code for single-image frames, $\mathcal{G}_{s}$, by training an Extractor CNN as its replacement.
-While $\mathcal{G}_{s}$ can be executed directly on the dataset, the processing pipeline is inherently non-batchable, as it handles image segmentations independently (see Appendix VII-B for examples).
-This non-batchability severely impacts RL training, causing slowdowns of 200–2000× compared to batchable operations (Extractor CNN), as demonstrated in Table III.
-From practical experience, this would translate to training times exceeding multiple weeks on an H40-level GPU for a single policy to converge, making this infeasible.
-To constrain computing time to a reasonable range, we train an Extractor CNN to replace $\mathcal{G}_{s}$ and efficiently compute $\mathcal{F}_{s}$, the features derived from a single image frame.
-The Extractor CNN is specifically designed to resolve the slowdown caused by segmentation-based analysis in $\mathcal{G}_{s}$.
-For each environment, we train an Extractor CNN on 7,000 images collected from diverse trajectories, validate it on 2,000 images, and test it on 1,000 images.
-The models achieved a mean squared error (MSE) consistently below 0.02 for each feature in the output vector when compared to the VLM-generated ground truth.
-Table II provides detailed performance statistics.
-These results confirm that the Extractor CNN is a reliable and computationally efficient replacement for $\mathcal{G}_{s}$.
-TABLE III: Runtime speedup for processing a batch of images across different environments on the same device (averaged over 10 iterations).
-Results denoted with * are derived from a different device with a better gpu.
-Env. Batch VLM (s) CNN (s) Speedup
-Skiing 8 0.870 0.011 81.02x
-16 1.745 0.003 523.37x
-32 3.523 0.005 743.77x
-64 7.055 0.008 877.90x
-128 14.449 0.021 693.74x
-Cart-pole 8 12.726 0.017 738.57x
-16 25.679 0.040 637.06x
-32 50.204 0.066 763.96x
-64 99.660 0.111 898.09x
-128* 254.484 0.097 2617.22x
-Boxing 8 0.400 0.017 23.07x
-16 0.785 0.004 191.03x
-32 1.581 0.007 238.51x
-64 3.146 0.009 344.34x
-128 6.240 0.025 254.40x
-IV-C Training Results
-Figure 3: Visualization of ICT in the Skiing Environment.
-Figure 4: Demonstration of the discrepancy between the physics simulation and the visualization.
+Details are included in Appendix A.2.
+4.2.3 Extractor module performance analysis
+In Step 3 of our methodology, we address the inefficiencies of the VLM-generated feature extraction code for single-image frames, $\mathcal{G}_{s}$, by training an lightweight Extractor Module as its replacement.
+While $\mathcal{G}_{s}$ can be executed directly on the dataset, the processing pipeline is inherently non-batchable, as it handles image segmentations independently.
+This non-batchability severely impacts RL training, causing slowdowns of around 20000× compared to batchable operations (our CNN Extractor Module), as demonstrated in Table 3.
+From practical experience, this would translate to training times exceeding multiple weeks on an H40-level GPU for a single policy to converge.
+We constrain computing time to a reasonable range with a lightweight Extractor Module, resolving the slowdown caused by segmentation-based analysis in $\mathcal{G}_{s}$.
+For each environment, we train an CNN on 7,000 images collected from diverse trajectories, validate it on 2,000 images, and test it on 1,000 images.
+Compared to the VLM-generated ground truth, our models achieve high accuracy (typically >0.95) on most binary features, while MSE remains low (<0.005) for the majority of continuous features, with one exception where error is notably higher.
+Table 1 provides detailed performance statistics.
+These results confirm that a lightweight vision model can serve as a reliable and computationally efficient replication for VLM generated extraction code $\mathcal{G}_{s}$.
+Table 1: Performance of the Extractor Module across environments, compared with ground truth from VLM-generated code.
+Continuous components are evaluated with MSE; boolean components are evaluated with accuracy.
+Dashes indicate vacant entries due to varying feature lengths and types across environments.
+Full feature descriptions are provided in Appendix A.2.2.
+Environment Feature Vector Indices (1–7)
+1 2 3 4 5 6 7
+Skiing (MSE) 0.00023 0.00025 0.0020 0.0039 0.0024 0.00034
+BabyAI (MSE) 0.0001 0.0001
+BabyAI (Acc) 1.000 0.979 0.975 0.972
+Highway (MSE) 0.0000 0.0000 1.153 0.0268 0.0209
+Highway (Acc) 0.988 0.948
+4.3 Training results
 We train ICT, our proposed interpretable and verifiable method that learns a tree structure with RL, alongside several baseline models across all environments.
@@
   • VLM-ICT (Rand Init): Our proposed method, but initialized randomly and not trained with RL.
-  • VLM-DT Heuristic: A baseline where VLM generates a heuristic tree to connect the semantic feature space with the action space.
   • VLM-MLP: An uninterpretable benchmark that connects a multi-layer perceptron between the semantic feature space and the action space.
   • Random Policy: A policy that selects actions by randomly sampling from the action space.
+(a) Atari Skiing
+(b) BabyAI: GoToRedBall
+(c) OpenAI Gym Highway
+Table 2: Runtime speedup on Skiing.
+Other domains in Appendix A.3.
+Batch VLM Code CNN Speedup
+8 19.056 0.001 13831.3x
+16 38.288 0.002 17325.6x
+32 76.212 0.004 19935.0x
+64 151.836 0.007 21692.9x
+128 308.468 0.014 22772.4x
+Figure 3: Results across three domains. Higher scores are better. Runtime comparison shows speedup of CNN-based extractor over VLM code.
+In addition to standard learning-based baselines, we also explore whether a Vision-Language Model (VLM) can directly perform action selection via prompting, given access to a semantic feature space it previously constructed.
+We prompt the VLM to synthesize an if-else style decision tree and obtain two variants of this baseline:
+  • VLM-DT (ChatGPT-4o-latest): A baseline where ChatGPT-4o generates a heuristic decision tree for action selection.
+  • VLM-DT (claude-3.5-sonnet): An equivalent baseline using Claude 3.5 Sonnet as the prompting engine.
+We use Proximal Policy Optimization (PPO) as the reinforcement learning algorithm to train our policy [24], with a learning rate of 0.0005 and a discount factor of 0.99.
+During testing, the policy is tested five times using different random seeds, each distinct from those used during training and evaluation.
+Results are shown in Fig.
+3.
 Across all three environments, VLM-ICT consistently outperforms other uninterpretable models (MLP) built on the semantic feature space, and demonstrates performance close to the CNN+RL black-box benchmark representing the upper limit of task performance.
-Notably, in the Cart-pole environment, VLM-ICT even outperforms the black-box CNN model.
-This is because the CNN fails to accurately capture the pole’s angle, a critical feature for formulating an effective policy.
 These consistent results highlight the robustness and effectiveness of our proposed method.
-The learned ICT in Skiing domain is shown in Fig.
-3.
-This tree was generated fully autonomously using the VLM $\rightarrow$ CNN $\rightarrow$ RL pipeline.
-The learned trees (ICT) are interpretable both subjectively and objectively, provided they remain small enough, as demonstrated in a previous user study [6].
-For the visualization of our learned ICT in Boxing and Cart-pole domains, please refer to Appendix VII-A.
-V Limitations and Future Works
-In the Cart-pole environment, we observe that the performance of trained VLM-ICT and VLM-MLP has yet to match that of RL policies trained using ground truth features [39, 6].
-This discrepancy stems from the gap between the underlying physics simulation and its visual representation, as illustrated in Fig.
-4.
-While the trained Extractor CNN’s predictions closely approximate the feature vectors extracted by the VLM-generated code, they still exhibit deviations from the ground truth feature vectors.
-This highlights a fundamental challenge in semantic RL from visual observations, where the extracted features may not perfectly align with the underlying physical state.
-Despite this limitation, our framework still outperforms the CNN-based policy, demonstrating the promise of VLMs in RL.
-To the best of our knowledge, we are the first to attempt solving the Cart-pole problem using image-based observations, demonstrating a novel and promising direction for interpretable RL.
-Future works could explore RL-based backpropagation to fine-tune the VLM models, enabling large language models to iteratively refine their outputs by learning from feedback generated through interactions with RL agents [40].
+Building on a semantically meaningful feature set, our ICTs contains nodes directly corresponding to human-understandable concepts.
+While we do not evaluate interpretability directly, this approach is consistent with prior work where performance serves as a proxy for meaningful representations [14, 36].
+A previous human study [18] also supports the interpretability of such trees.
+We visualize the learned ICT in Appendix A.1, along with example cases demonstrating how the ICT yields plausible local explanations in example scenarios.
+5 Limitations and future works
+Our framework is yet limited by the assumption that SAM performs reasonably well across environments, which may hinder generalizability in domains with poor segmentation.
+Future work could explore RL-based backpropagation to fine-tune VLMs, allowing them to iteratively improve outputs based on agent feedback [1].
 This approach has the potential to enable VLMs to identify discrepancies between the underlying physics simulation and its visual representation, leading to improved feature extraction for downstream RL tasks.
-VI Conclusion
-We introduce SILVA, a novel framework for achieving automated semantic interpretability in RL.
+6 Conclusion
+We introduce iTRACE, a novel framework for achieving automated semantic interpretability in RL.
 It automates feature space construction and extraction with zero-shot VLM querying, making it generalizable to unseen environments without human-in-the-loop.
-To overcome VLMs’ inability to generate valid decision policies, SILVA leverages an ICT for RL training, obtaining transparent, human-understandable, and verifiable decision rules.
-The framework outperforms black-box models built on the same feature space and achieves performance on par with purely uninterpretable black-box models.
-While all modern interpretable models, including tree-based approaches, face challenges in complex environments, SILVA stands out among them as the first to achieve automated semantic interpretability in RL, marking a significant advancement that underscores its impact.
-Ethics and Impact Statement
-SILVA enhances interpretability in RL by automating semantic feature extraction via VLMs, reducing reliance on human annotations and fostering transparent AI decision-making.
-By structuring policies in an interpretable tree-based format, SILVA aligns with ethical AI principles emphasizing explainability, accountability, and trust.
-However, reliance on VLMs introduces biases from pre-trained models, potentially causing discriminatory behavior or unreliable feature extraction.
-While SILVA improves interpretability, it still depends on black-box models for feature generation, challenging full transparency.
-To mitigate these risks, future work could incorporate bias audits, human-in-the-loop verification, and standardized fairness metrics.
+To overcome VLMs’ inability to generate valid decision policies, iTRACE leverages an ICT for RL training, obtaining transparent, human-understandable, and verifiable decision rules.
+The framework achieves performance on par with purely uninterpretable black-box models and black-box models built on the same feature space.
+While all modern interpretable models, including tree-based approaches, face challenges in complex environments, iTRACE stands out among them as the first to achieve automated semantic interpretability in RL, marking a significant advancement that underscores its impact.
+Appendix A Appendix
+A.1 Appendix A: ICT Visualization
+Figure 4: Visualization of a highlighted decision path (outlined in red) in the BabyAI GoToRedBall environment, with the corresponding agent observation shown on the right.
+In this scenario, the agent’s state vector indicates it is close to the left wall (Left Wall = 1), not close to the front or right walls (Front Wall = 0, Right Wall = 0), and an object is detected (Obj Exist = 1) at ego-centric coordinates (Obj_x = 1, Obj_y = 4).
+At this sceneario, our learned ICT suggest the agent to move forward which aligns with humans intuition.
+Figure 5: Illustration of a highlighted decision path (outlined in red) in the BabyAI GoToRedBall environment.
+The corresponding agent observation is shown on the right.
+In this scenario, the agent’s state vector indicates it is not close to any wall (Left Wall = 0, Front Wall = 0, Right Wall = 0), and no object is detected (Obj Exist = 0).
+The default object location is set to Obj_x = 0, Obj_y = –5.
+Since the object is not presented in agent’s view, our learned ICT suggests the agent to turn left to search for the object.
+A.2 Appendix B: VLM Prompting Ablation
+A.2.1 Segmentation Labeling
+Among all evaluated environments, Skiing posed the greatest challenge for LLM-based segmentation labeling.
+To reduce expensive API usage while showcasing a representative case, we center our analysis on this environment.
+As shown in Figure 6, segmentation quality is imperfect: the skier is sometimes only partially labeled or missed entirely due to large appearance variations across frames.
+BabyAI and Highway environments contain many repetitive segments that occupy a large portion of each image.
+To stay within the VLM’s context window token limit, we filter out strictly identical segments using simple heuristics such as aspect ratio and area.
+This assumes that highly frequent, repetitive elements are less critical for feature extraction.
+This limitation can be mitigated as LLM context capacity increases.
+Once longer context windows are supported, all segments can be included, and irrelevant decorative elements can still be correctly labeled and filtered, as shown qualitatively in our examples.
+(a) Actual skier is labeled correctly, but decoration elements at the bottom are labeled as partial skier.
+(b) Noisy segment is identified as skier, while the actual skier is labeled partial.
+(c) Skier identified as decoration.
+Figure 6: Examples of imperfect segmentations in the Skiing environment.
+Despite errors in example segmentation, our method successfully retrieves effective extraction functions for key features across environments, and the resulting code generalizes well across frames.
+This demonstrates that our pipeline is robust to segmentation errors: as long as a subset of segments is labeled correctly, the system can proceed to generate reliable feature extraction logic.
+Reproducibility.
+To support reproducibility, we provide the specific examples and their LLM segmentations for all three environments. Some of them are imperfect.
+(a) Skiing Example 1.
+(b) Skiing Example 2.
+(c) Skiing Example 3.
+Figure 7: Randomly selected segmentation examples we provide to the VLM.
+Example 1 is good while 2 and 3 are imperfect.
+Example 2 identifies the skier as a partial skier and a noisy segment as skier.
+Example 3 identifies both the actual skier on the right and decoration elements at the bottom as partial skier.
+(a) BabyAI example 1.
+(b) BabyAI example 2.
+(c) BabyAI example 3.
+Figure 8: Randomly selected segmentation examples we provide to the VLM. Some segmentation is imperfect. For example, the green box is not correctly identified in example 2.
+(a) Highway example 1.
+(b) Highway example 2.
+(c) Highway example 3.
+Figure 9: Randomly selected segmentation examples we provide to the VLM for Highway. Segmentation for this environment is mostly perfect.
+A.2.2 Feature Dictionary Extraction
+Here we provide the complete list of features that the VLM decided for each environment in Table. 3, 4, 5 .
+Table 3: Feature schema for BabyAI RedBall environment.
+Object Feature Type Requires Temporal Context
+red_ball x_position Continuous No
+y_position Continuous No
+is_visible Categorical No
+obstacle_forward is_blocked Categorical No
+obstacle_left is_blocked Categorical No
+obstacle_right is_blocked Categorical No
+Table 4: Feature schema for Highway environment.
+Object Feature Type Requires Temporal Context
+agent x_position Continuous No
+y_position Continuous No
+x_velocity Continuous Yes
+lane_existence upper_lane_exists Ordinal No
+lower_lane_exists Ordinal No
+vehicle_ahead_same_lane x_position Continuous No
+x_velocity Continuous Yes
+vehicle_adjacent_upper_lane x_position Continuous No
+x_velocity Continuous Yes
+vehicle_adjacent_lower_lane x_position Continuous No
+x_velocity Continuous Yes
+Table 5: Feature schema for Skiing environment.
+Object Feature Type Requires Temporal Context
+next_gate x_position_left_pole Continuous No
+x_position_right_pole Continuous No
+y_position Continuous No
+closest_tree x_position Continuous No
+y_position Continuous No
+skier x_position Continuous No
+x_velocity Continuous Yes
+A.2.3 Extraction Function Retrieval
+Table 6: Prompting statistics for VLM-generated function 1 across 50 trials.
+“Code failure” indicates code that cannot import or run.
+“Excluded as constant" refers to code that generates invariant outputs across diverse samples and is therefore excluded.
+“Majority selection coverage” represents the proportion of code yielding identical outputs in the majority voting process.
+Step 1 refers to the process of generating $\mathcal{G}_{s}$, while Step 2 refers to the process of generating $\mathcal{G}_{m}$.
+Env. Code Failure Rate (%) Excluded as Constant (%) Majority Selection Coverage (%)
+Skiing 8 0 72
+BabyAI 8 74 4
+Highway 0 90 4
+To address the unreliability of VLM-generated code, particularly hallucinations and execution failures, we implement a selection procedure to identify robust feature extraction functions.
+For each of the three environments, we prompt the VLM to generate 50 candidate functions, evaluate them on a dataset of 1000 images, and filter out those that either fail to run or produce invariant outputs across all samples.
+From the remaining valid functions, we apply majority voting based on output similarity and select the function whose outputs most frequently align with others.
+This strategy assumes that the dataset is sufficiently diverse such that no single feature remains constant across all frames.
+Table 6 presents statistics from this process, including code failure rates, exclusions due to constant outputs, and majority selection coverage.
+While the majority selection coverage appears relatively low, this is primarily due to inconsistent normalization of features representing relative distances.
+In contrast, features encoding object positions tend to be represented more consistently across functions.
+A.3 Appendix C: CNN Performance Analysis
+A.3.1 CNN structure for the Skiing environment.
+The SkiingCNN model is a general-purpose convolutional network designed for multiple environments.
+For the skiing environment specifically, it outputs a 6-dimensional feature vector per frame, capturing semantic information such as skier position, gate pole positions, and nearby trees.
+  • Input: A $3\times 56\times 76$ RGB segmentation image, with distinct segments corresponding to skier, gates, and trees.
+  • Environment Routing: At initialization, the network identifies its task from an environment string (e.g., env_n = "skiing").
+This string determines the output dimensionality; for skiing, it is fixed to 6.
+The model uses a compact, task-specific CNN with the following layers:
+  • Conv2d(3 $\rightarrow$ 32, kernel=5) $\rightarrow$ ReLU $\rightarrow$ MaxPool2d(2) $\rightarrow$ [32, 25, 19]
+  • Conv2d(32 $\rightarrow$ 64, kernel=3) $\rightarrow$ ReLU $\rightarrow$ MaxPool2d(2) $\rightarrow$ [64, 5, 4]
+  • Conv2d(64 $\rightarrow$ 64, kernel=3) $\rightarrow$ ReLU $\rightarrow$ [64, 3, 2]
+This yields a low-resolution spatial encoding with 64 channels and a flattened size of 384, computed dynamically during initialization.
+MLP Output Head. The fully connected layers process the flattened feature vector to produce semantic outputs:
+  • Linear(flat_size $\rightarrow$ 512) $\rightarrow$ ReLU $\rightarrow$ Dropout(0.4)
+  • Linear(512 $\rightarrow$ 6) for skiing outputs
+Output Format.
+The final 6-dimensional output contains semantically meaningful quantities related to skiing dynamics.
+Unlike in some other environments (e.g., RedBall), no post-hoc sigmoid activation is applied, as all skiing outputs are treated as continuous and directly regressed.
+A.3.2 CNN structure for the BabyAI-GoToRedBall environment.
+For this part, we use a combination of two CNN networks, BinaryDetectCNN and BallDetectionCNN.
+The BinaryDetectCNN is a multi-task convolutional model designed to predict (1) binary obstacle existence in the forward, left, and right directions, and (2) the existence of a red ball in the frame.
+The model uses shared convolutional features and two dedicated branches:
+  • Shared Feature Extractor: Three convolutional layers with channels [3 $\rightarrow$ 32 $\rightarrow$ 64 $\rightarrow$ 128], each followed by batch normalization and ReLU.
+The feature map is downsampled with stride-2 convolutions instead of pooling for better spatial preservation.
+  • Block Detection Head: A dedicated encoder (1 conv layer) followed by adaptive average pooling and a two-layer MLP [128 $\rightarrow$ 256 $\rightarrow$ 3] for predicting obstacle presence in three directions.
+The output is passed through sigmoid activation.
+  • Ball Existence Head: A parallel encoder with the same structure, yielding a single sigmoid output indicating the presence of a red ball.
+  • Output Format: The final output is a 6-dimensional vector: 3 binary obstacle indicators, 2 position placeholders, and 1 binary ball existence score.
+The placeholder values are used to maintain a consistent output shape and are overwritten when positional regression is available.
+The BallDetectionCNN is a separate network responsible for regressing the $(x,y)$ position of the red ball, but only if the ball is detected by BinaryDetectCNN.
+  • Convolutional Backbone: A four-block CNN with increasing channels [3 $\rightarrow$ 32 $\rightarrow$ 64 $\rightarrow$ 128 $\rightarrow$ 256], each block consisting of convolution, batch normalization, ReLU, and $2\times 2$ max pooling.
+  • Adaptive Pooling: The output is pooled to a $1\times 1$ spatial map to support arbitrary input sizes.
+  • Regression Head: A three-layer MLP [256 $\rightarrow$ 128 $\rightarrow$ 64 $\rightarrow$ 2] outputs the normalized coordinates of the ball.
+  • Inference Logic: The regressor is only invoked if the ball existence score from BinaryDetectCNN exceeds 0.5.
+The image is resized to $64\times 64$ and normalized before being passed to the regressor.
+At inference time, both models are used jointly to generate a compact 6-dimensional feature vector:
+[left_blocked, front_blocked, right_blocked, ball_exists, ball_x, ball_y]
+If the ball is not detected, the position fields are filled with default placeholder values (e.g., $x=0$, $y=-5$).
+This conditional two-stage pipeline enables efficient and accurate feature extraction without overfitting to spatial priors or dense supervision.
+A.3.3 CNN structure for the Highway environment.
+To improve regression accuracy for specific vehicle positions, we deploy three additional networks, each a specialized regressor based on ResNet-18. These models handle:
+  1. 1.
+    Vehicle ahead in the same lane
+  2. 2.
+    Vehicle in the adjacent upper lane
+  3. 3.
+    Vehicle in the adjacent lower lane
+Each HighwayCNN model follows the structure below:
+  • Backbone: A ResNet-18 pretrained on ImageNet. All layers are initially frozen, with the last three residual blocks selectively unfrozen for fine-tuning.
+  • Custom Head: The original fully connected layer is replaced with a regression head: [512 $\rightarrow$ 256 $\rightarrow$ 128 $\rightarrow$ 1], with batch normalization, ReLU activations, and dropout (0.3, 0.2).
+  • Output: A single scalar value representing the $x$-position of the target vehicle.
+At inference time, the models are used jointly to construct a 7-dimensional feature vector per frame:
+[lower_lane, upper_lane, agent_x, agent_y, vehicle_ahead_x, vehicle_upper_x, vehicle_lower_x]
+The main CNN provides lane existence and agent position.
+The three auxiliary regressors are queried independently to estimate the x-coordinates of nearby vehicles in the ego, upper, and lower lanes.
+This multi-stage pipeline allows for flexible specialization across sub-tasks while maintaining a compact, semantically meaningful feature representation.
+A.3.4 Additional Runtime Analysis
+We used a more complicated CNN structure which slows down CNN exponentially in those two domains, but the CNN computational time is still significantly non-trivially shorter than the actual computing time.
+Table 7: Runtime speedup on BabyAI
+Batch VLM Code CNN Speedup
+8 42.435 (s) 0.037 (s) 1132.93x
+16 97.351 (s) 0.042 (s) 2303.66x
+32 187.182 (s) 0.097 (s) 1924.26x
+64 376.923 (s) 0.112 (s) 3361.79x
+128 750.582 (s) 0.342 (s) 2196.65x
+Table 8: Runtime speedup on Highway.
+Batch VLM Code CNN Speedup
+8 42.421 (s) 0.025 (s) 1683.08x
+16 88.326 (s) 0.040 (s) 2207.01x
+32 172.973 (s) 0.084 (s) 2047.23x
+64 347.060 (s) 0.186 (s) 1868.32x
+128 696.099 (s) 0.328 (s) 2122.57x
+A.4 Appendix D: Prompts
+We employ Chain-of-Thought prompting combined with structurally formatted Markdown outputs to guide the model toward producing target results.
+The complete prompts used for segmentation labeling, feature extraction, Function 1 generation, and Function 2 generation are provided below.
+A.4.1 Segmentation Prompt
+You are tasked with interpreting every segment from a segmented image and labeling it with what it appears to represent.
+You will be given a single image called image.png, along with a segmentation file masks.json generated by Segment Anything.
+This file contains a list of all visual regions detected in the image.
+Each segment includes attributes such as bounding box, area, dominant color, and other visual metadata.
+Your goal is to assign a predicted_object label to every segment in this list, whether the segment depicts a featured object, a decorative element, or a fragment of a larger object.
+Note that the color attribute is given by the most dominant color in the segment. If the segment is colorful, the dominant is not reliable and you should not depend on it.
+You will first receive a feature dictionary specifying which features are to be derived from the image.
+Begin by examining this dictionary to determine which object types are necessary in order to compute the values of those features.
+You may ignore any features marked with "requires_temporal_context": True, as they are not relevant for single-frame analysis.
+Based on the remaining features, deduce what kinds of objects must be identified to extract those values.
+These are your featured objects, and you must list them explicitly before proceeding to segment analysis.
+The feature dictionary is defined here: <insert-features>
+Domain Knowledge: <insert-domain-knowledge>
+Do not assume that visually large, centered, or high-contrast segments are featured objects unless it aligns with what you see from the image I provided.
+A segment should only be labeled as a featured object if its shape, size, position, and color closely match a plausible instance of that object in the image.
+The presence or prominence of an object in the image is not evidence of its relevance unless it is explicitly needed for feature extraction.
+If the feature dictionary requires you to distinguish objects by their relative positions (e.g., “ahead”, “behind”, “left of”, “right of”, or “same group”), you must assign them a general segmentation label, and leave any spatial analysis alone.
+A separate stage will handle positional analysis later, so you don’t need to worry about it here.
+Once you have identified the featured objects, begin a two-pass labeling process over the segmentation data.
+For the first pass, analyze each segment independently.
+Use only the attributes provided in the masks.json file—such as bounding box, area, aspect ratio, dominant color, and relative position—and your visual observation of the image.
+Base your predictions strictly on observed features and metadata.
+Do not infer identities from prior knowledge or gameplay expectations.
+Assign each segment a predicted_object label that reflects what it is most likely to represent.
+If a segment matches one of the featured objects you previously identified, label it with that object’s name.
+If a segment appears to be a visual decoration (such as background, UI overlay, or border), label it explicitly as such.
+If a segment clearly represents only a part of a larger object, assign it a label indicating it is a fragment (e.g., "fragment" or "partial_<object>").
+If you cannot confidently classify a segment, label it as "ambiguous".
+After this first pass, review your labeling decisions in context.
+Consider whether your predictions make sense collectively.
+For example, ask whether any object that should only appear once is represented by multiple segments, whether any key object is missing entirely, or whether the traits of any labeled segments contradict the visual evidence from the image.
+If you identify inconsistencies, refine your labels accordingly to make the overall scene interpretation more coherent.
+The second pass should resolve overlaps, mismatches, or missing predictions that become evident only in context.
+Your output must return the list of segments from the original masks.json, preserving all original fields while adding a new "predicted_object" field to each segment dictionary.
+Every segment must be labeled.
+No segment may be left without a prediction, even if it is decorative, irrelevant, or ambiguous.
+Segmentation data derived from the image is here: <insert-segment-knowledge>
+Begin your reasoning inside the Reasoning section, where you strictly follow the thinking process:
+  1. 1.
+    Ignore the segmentation and look at the image. Describe all elements you observe, including all decorational ones.
+  2. 2.
+    Look at the feature dictionary. Deduce the featured objects required by the feature dictionary.
+  3. 3.
+    Analyze each segment independently and assign them with a prediction label following the instructions for the first pass.
+  4. 4.
+    Conduct the second pass.
+Finally, return the updated segment list with predicted_object annotations inside the ## Final Answer section.
+Your final answer should follow the same format with the input, strictly copying the values while adding your predictions:
+    "area": area,
+    "bbox": [x, y, w, h],
+    "predicted_iou": predicted_iou,
+    "point_coords": [
+      [x, y],
+    "stability_score": stability_score,
+    "crop_box": [x, y, w, h],
+    "dominant_color": [r, g, b],
+    "predicted_object": object_name
+Your output should follow this structure:
+## Reasoning
+…all steps…
+## Final Answer
+  { "area": ..., "predicted_object": "your prediction", ... },
+A.4.2 Feature Extraction Prompt
+Instructions: You are given a series of image frames from a specific environment, and your task is to determine the features essential for making the immediate decision.
+Focus on extracting the positional data of key objects and any associated features that can aid in choosing the immediate movement.
+Ignore any elements that are visible but do not affect the agent’s ability to choose its immediate movement.
+This includes elements whose behavior is entirely deterministic once triggered or which cannot be altered by future actions.
+You should also ignore anything that cannot be directly inferred from the image or requires assumptions beyond visual evidence.
+If certain properties of an object (such as its type, color, or exact shape) do not influence what the agent will do next, you are encouraged to abstract them away.
+The goal is to capture only the information needed to distinguish between available actions.
+Extraneous visual or semantic detail should be excluded, even if it is visible in the frame.
+You may also consider changes in object positions across time (e.g., motion trends) when necessary.
+In some environments, both horizontal and vertical positions/velocities of key elements across frames are informative.
+They should be always separated by direction and axis (e.g., x_position, y_position, x_velocity, y_velocity).
+Positional or motion features must be separated into explicit horizontal (x) and vertical (y) components whenever relevant.
+If either direction (horizontal/vertical) doesn’t matter, exclude it; otherwise, always include both horizontal and vertical direction.
+If movement trends are not relevant to decision-making in a given case, they should be omitted entirely.
+Not all information visible in the image is essential for decision-making. Disregard anything that isn’t.
+Important: Some environments may contain multiple instances of the same type of object.
+You should not assume a fixed or singular representation for such objects unless explicitly outlined in the environment description.
+When considering non-goal objects that could impact the movement (obstacles, enemies, etc.), avoid modeling the entire map or distant entities.
+Instead, identify whether anything in the agent’s immediately reachable space would prevent it from either choosing or executing any of its primitive actions.
+Most commonly, this will be the closest instance (obstacles, enemies, etc.), as that is most likely to impact immediate behavior.
+You must not include other instances unless their effect is simultaneous and unavoidable.
+You should also explain your selection criterion under Final Description.
+When identifying environmental elements that constrain the agent’s immediate behavior, consider whether anything in the agent’s directly actionable space—that is, any region reachable by a single primitive action—would limit or prevent the choice of that action.
+For instance, an obstacle to the side may not block movement directly but could influence whether turning should be chosen as the next immediate action.
+Do not include distant or visually salient elements unless their effect is immediate and unavoidable.
+You should also abstract away any properties (such as type, color, or identity) that do not influence the outcome of the agent’s next action.
+You should not ground your analysis in the specific frames you’re shown, except if things stay consistent across them.
+These are only examples to help illustrate the structure of possible observations.
+Your job is not to analyze the example frames individually, but to construct a feature dictionary that would apply across frames and across environments sharing the same interaction mechanism.
+This means the actual position or visibility of the target object in the examples should not influence your output.
+Domain knowledge: <insert-domain-knowledge>
+Your answer should be structured as:
+  "element 1": [
+        "attribute 1": {
+            "values": ["option1", "option2", ...],
+            "type": "ordinal",
+            "requires_temporal_context": False
+        "attribute 2": {
+            "values": x,
+            "type": "continuous",
+            "requires_temporal_context": False
+        "attribute 1": {
+            "values": x,
+            "type": "continuous",
+            "requires_temporal_context": True
+        "attribute 2": {
+            "values": x,
+            "type": "continuous",
+            "requires_temporal_context": False
+  "element 2": {
+        "attribute 1": {
+            "values": ["option1", "option2", ...],
+            "type": "ordinal",
+            "requires_temporal_context": True
+Each element should represent an environmental object or factor that could influence the agent’s immediate decision.
+If multiple elements of the same type are relevant, they should be provided as a list.
+Each attribute describes a property of that element that could affect the decision-making process.
+Please do not make up or fill in specific values like 60, 100, or -5 in the "values" field.
+That field should remain blank or placeholder-like (e.g., "values": "continuous") to indicate the type and temporal context of each attribute.
+Your output should follow this structure:
+## Reasoning
+…all steps…
+## Final Answer
+(stricly follow the format specified above)
+## Description
+Describe the visual traits of the featured objects and the selection mechanism from a list of identical objects (if applicable).
+This should help a third-party reader understand what your features refer to, not explain your reasoning process.
+A.4.3 Function 1 Generation Prompt
+You are tasked to write a python function to retrieve values of the specified features from an image in a consistent environment.
+This function should be able to generalize across frames.
+Describe all positions by ratio in the image x y axis.
+Your code should take in one single image frame (image.png) along with a json dictionary (masks.json) given by Segment Anything on the image.
+Name the function you write ‘extract_features(img_dir, json_dir)‘.
+Note that the segments in the JSON dictionary contain every possible element in the image frame, many of which are irrelevant or decorative.
+Thus, you must include logic to exclude such elements and only retain the segments that are visually meaningful for the defined features.
+Each entry in masks.json corresponds to a visually distinct segment in the image.
+Note that these segments include both relevant objects (e.g.
+boxers) and irrelevant or decorative elements (e.g.
+background, UI, borders, large static surfaces, or visually dominant non-interactive areas).
+Note that this segmentation file provides a list instead of a dictionary.
+I will provide you with some example images and their segmentation data printed in text form, but note that the function you write will deal with a completely different image and its different segmentation data.
+The example images are embedded with visualizations of the segmentations to give an idea of what the targets and the noises will look like, but the actual image your function deals with won’t be.
+In the example segmentation data, each segment includes a "predicted_object" field which labels what the segment likely represents.
+This field exists only for the sake of demonstration—to help you learn what kinds of attributes distinguish key objects from background clutter.
+The "predicted_object" field will NOT be present in the actual masks.json file that your function processes.
+Therefore, your final code must NOT rely on this field in any way.
+Instead, use it in the examples purely as reference to derive the selection logic.
+However, you must study the "predicted_object" labels in the examples carefully to identify consistent, distinguishing traits for each object type you intend to detect.
+Your selection logic should be explicitly based on those shared traits—such as area, color composition, aspect ratio, and position—that reliably differentiate the intended objects from irrelevant segments like UI, background, or decorative fragments.
+Do not use overly broad filters unless you have confirmed—through all provided examples—that such filters exclude all non-target segments without failure.
+Important: Raw segmentation masks are not provided. You must instead analyze the image directly using the information from the image pixels as a visual clue if needed.
+Some features may require information that goes beyond what’s available in the segments alone.
+While segments provide a reliable reference for object-level statistics, they may not capture everything—such as irregular shapes that aren’t well represented by bounding boxes or aspect ratios, or subtle lighting conditions where average color is insufficient.
+In such cases, you should refer directly to the image rather than relying solely on the segmentation.
+Segmentation data derived from the example images: <insert-segment-knowledge>
+You should determine the feature objects you want to search for based on the domain knowledge and target feature dictionary I give you, and think about how to select them from the segmentation data.
+Domain knowledge: <insert-domain-knowledge>
+Features that the function is tasked to derive: <insert-features>
+where each individual dictionary item represents an environmental object or factor that could influence the agent’s next decision.
+If you identified multiple objects of the same type, they should be provided as a list in the dictionary item as specified by examples in the given feature dictionary.
+For the attributes of each object, there will be a "requires_temporal_context" field which is either True or False (boolean).
+If it’s False, neglect the attribute and don’t return anything.
+Only if it’s True, analyze the image and the Segment Anything masks to decide a value.
+For the attributes of each object, there will be a "type" field which would be either "continuous" or "ordinal".
+If marked "continuous", your code should derive a number value.
+If marked "ordinal", your code should return one of the specified ordinal values as a string.
+Additional clarifications about the features can be found here.
+You should use it either (1) as supplementary information about the object’s visual traits to help your selection, or (2) as strict criteria that must be enforced when selecting a type of objects, if any specific requirements are stated.
+<insert-explanation>
+Your code should return a dictionary keeping all the original dictionary items, while only including attribute values and excluding the "type" and "requires_temporal_context" fields.
+Your code should be able to run as it is, without requiring further information or edit.
+Note that every featured object exists in the example image, but not necessarily in all images the function deals with, so return None if you can’t find the featured object.
+Segment Anything already provides well-grouped regions for each object — do not merge or cluster multiple segments.
+If the featured object is supposed to appear only once, you must select exactly one segment to represent it.
+Note that the color attribute is given by the most dominant color in the segment. If the segment is colorful, the dominant is not reliable and you should not depend on it.
+Objects will vary across frames.
+You are exposed to only a limited sample of image frames, so it’s advised to keep the selection criteria loose as long as you can verify it doesn’t lead to misselecting any decorative elements, UI backgrounds or fragments.
+If the selection criteria involves spatial relationships and occupancy between general objects, you must NOT use the object’s bounding box or center to make this decision.
+Instead, you should use pixel-wise content in the image region corresponding to the spatial location and determine the featured value based on its actual pixel content.
+Regions that are not occupied tend to exhibit consistent visual characteristics—such as uniform texture or stable color patterns—while occupied regions vary more widely in appearance.
+You should resort to the domain knowledge for additional information and visual cues of representative patterns observed in known examples.
+If the selection criteria involves spatial relationships between fully-filled objects, you should account for the object’s full bounding box, not just its center.
+However, this is only allowed when the goal is to estimate a continuous relative offset, not binary occupancy.
+An object may influence or obstruct another’s path if any part of its area extends into the relevant region, even if its center does not.
+If the feature dictionary involves discrete categories of relative positions, you must NOT divide the image into uniform bands or grids along the height, as the viewpoint may shift across frames.
+Instead, you should define fixed-width zones based on relative spatial relationships of visual landmarks or reference objects visible in the current frame.
+Additional Instruction: If a featured object is represented by multiple segments of the same type, it is acceptable—and sometimes necessary—to first retrieve a related value that is not included in the final feature dictionary to aid in the selection process.
+Strictly follow the defined selection criteria, and do not hesitate to compute temporary or auxiliary values if they are essential for choosing the correct segment.
+However, these intermediate values should not be included in the final dictionary output unless they are explicitly required.
+Your thinking should follow these steps.
+– Look at the segmentations for each example image and the "predicted_object" field (only available in the examples), assuming that the prediction is correct.
+Pay attention to the attributes of featured objects – what traits can be used to select them from the segmentation json?
+– Discuss and outline your approach for separating decorative elements from the specified key objects.
+Would any of the decorative elements, especially borders, UI, background, fall into your selection category?
+How did you ensure that the segment you selected is the actual item?
+If the item is supposed to be singular instead of a list, how did you ensure only one segmentation would qualify under your criteria as the actual item?
+– Describe your object selection logic and your plan of writing a python function that detects the positions of the objects of interest detailedly.
+If some segmentations are clearly decorative or part of the background (e.g., one of their statistics is noticeably different), it’s recommended to filter them out before selecting the actual object to avoid errors.
+Again, since objects will vary across frames, you should keep the object selection criteria as loose as possible.
+– Simulate your planned selection logic on the full segmentation data, one by one, including not just featured objects but also irrelevant segments, UI backgrounds, and partial fragments.
+You must verify that your method consistently identifies the correct segments and does not mistakenly include irrelevant ones.
+Focus on the numerical values on the first segmentation example I gave you.
+If your strategy fails (e.g., selects a background or UI element), refine it before proceeding.
+You must prove to yourself that the approach works on the provided examples.
+You must validate that your approach works without relying on "predicted_object" in actual usage.
+## Reasoning
+## Code
+A.4.4 Function 2 Generation Prompt
+You are tasked to write a python function to retrieve values of the specified features from two chronological images in a consistent environment, where object traits remain stable.
+The timestamp between the images will always stay consistent.
+Your code should take in two image frames (png) along with their associated json dictionaries (json) given by Segment Anything on the images.
+Your function should be named after ‘def extract_features_two_frames(img1_dir, json1_dir, img2_dir, json2_dir)‘.
+Domain knowledge: <insert-domain-knowledge>
+Features that the function is tasked to derive: <insert-features>
+where each individual dictionary item represents an environmental object or factor that could influence the agent’s next decision.
+If you identified multiple objects of the same type, they should be provided as a list in the dictionary item as specified by examples in the given feature dictionary.
+For the attributes of each object, there will be a "type" field which would be either "continuous" or "ordinal".
+If marked "continuous", your code should derive a number value.
+If marked "ordinal", your code should return one of the specified ordinal values as a string.
+For the attributes of each object, there will be a "requires_temporal_context" field which is either True or False (boolean).
+This field indicates whether the attribute value requires two images frames or one single image frame to determine.
+In a previous stage, you already wrote a python function that takes in one single image frame and its associated segmentation information (input: image.png, masks.json) and returns a dictionary keeping all the original dictionary items, only including attribute values and excluding the "type" and "requires_temporal_context" fields.
+For each attribute, the function determines an actual value if "requires_temporal_context" is False, and inserts None if it’s True.
+The function is:
+<insert-function1>
+In this step, you should write a python code that utilizes the previous function for single image.
+Your code should take in two image frames (image1.png, image2.png) along with their associated json dictionaries (masks1.json, masks2.json), apply the single image function to them separately, and then return a dictionary keeping all the original dictionary items, only including attribute values and excluding the "type" and "requires_temporal_context" fields.
+Your code should be able to run as it is, without requiring further information or edit.
+Note that every featured object exists in the example image, but not necessarily in all images the function deals with, so account for that in your code logic.
+For each attribute, your code should return the value of running the previous function on the second image and completely ignore the value from the first image if "requires_temporal_context" is False.
+If "requires_temporal_context" is True, you should either calculate from the results of running the previous function on both images, or if that’s still not sufficient for the attribute, analyze the image and segmentation information from both images to retrieve a value.
+If a velocity or any other attribute requires using two values, and either value is None, then your code must return None for that attribute.
+Describe all positions by ratio in the image x y axis.
+I will provide you with three example images and their segmentation data printed in text form, but note that the function you write will deal with different images and different segmentation data.
+Segmentation data derived from the example images: <insert-segment-knowledge>
+Begin by explaining your reasoning step-by-step.
+After that, write the required python function.
+Double check that if an attribute requires using two values, either value being None would not affect your output.
+## Reasoning
+## Code
+A.4.5 Domain knowledge.
+The domain knowledge of each environment can be found below respectively.
+Skiing
+You control a skier who moves strictly downhill and can only shift sideways (left or right) by tilting your skis in that direction.
+The game uses a relative camera view: the skier remains approximately fixed near the top-center of the image (though not at the exact top), while the camera scrolls with the skier’s downhill motion.
+However, the camera does not follow the skier’s lateral movement, so their sideways position within the frame varies.
+In image space, the downhill direction corresponds to moving toward the bottom of the screen.
+The skier cannot move upward, so you must only consider gates and trees that appear below the skier in the image.
+The skier is primarily moving downhills towards the bottom of the image, which is also its primary goal.
+You must ignore gates and trees above the skier since they are no longer reachable.
+The goal is to pass through all gates as quickly as possible while avoiding trees below the skier.
+A gate consists of the horizontal space between two poles positioned at the same latitude.
+The skier must pass between both poles of the same gate to successfully clear it.
+You are penalized five seconds for each gate you miss.
+If you hit a gate or a tree, your skier will jump back up briefly before continuing downhill.
+You should avoid extreme sideways movement which is indicated by a noticeable drop in skier speed.
+In such situations, you must actively tilt the skis to regain speed and keep moving.
+Otherwise, the skier will gradually get stuck.
+Actions Skiing has the action space of Discrete(3) with the table below listing the meaning of each action’s meanings.
+Value | Meaning
+    0 | NOOP
+    1 | RIGHT
+    2 | LEFT
+Highway
+You control a car that drives forward along a multi-lane highway and can change lanes or adjust its speed to navigate traffic.
+The lanes appear as vertical bands (spanning top to bottom), and vehicles move horizontally across them from left to right.
+The road contains multiple parallel lanes arranged vertically, with some lanes located closer to the upper edge and others closer to the lower edge.
+Vehicles travel along these lanes and may appear in the same, adjacent upper, or adjacent lower lanes relative to the agent.
+Not every lane is directly next to an edge, and the number of visible lanes will vary depending on your position.
+Your task is to drive efficiently by anticipating traffic behavior and adjusting your actions to preserve forward momentum.
+You can switch lanes upward or downward, as long as such a lane exists and is not blocked.
+Other cars on the road may accelerate, decelerate, or change lanes, making them dynamic obstacles.
+Adjacent or closely ahead vehicles in the immediate upper or lower lanes present the most immediate risk during a lane change and should be closely monitored for lateral overlap and motion trends.
+The road is defined by visible lane markers: dashed lines separate adjacent lanes, while solid lines mark the outer boundaries.
+If the vehicle is next to a solid line, it means there is no adjacent lane available for switching in that direction.
+Veering beyond the highway boundaries would result in an immediate crash.
+The distance between lanes is uniform, meaning the gap between any two neighboring lane markers represents the standard lane interval.
+The width (height) of a car is notably smaller than the width (height) of an entire lane, and thus should not be used as a measure of lane width.
+The vertical distance between lane markers reliably measures lane width.
+Lane marker distance can be confidently used to determine if a car is in an adjacent lane or in further lanes.
+If you want lane width, you must calculate it as the distance between lane markers. You must NOT assert it from the segmentations.
+Collisions result in episode termination and must be avoided.
+You are rewarded for reaching and maintaining higher speeds without incident.
+Smooth driving and intelligent maneuvering through dense traffic are key to success in this environment.
+Actions Driving has the action space of Discrete(5) with the table below listing each action’s meaning.
+Value | Meaning
+0 | IDLE
+1 | FASTER
+2 | SLOWER
+3 | SWITCH_UPPER
+4 | SWITCH_LOWER
+BabyAI
+In this BabyAI GoToRedBall environment, the world is an enclosed grid.
+The agent is dropped into a new configuration each episode, surrounded by a few simple objects of different colors.
+Your goal is to go to the red ball.
+A complete list of all possible items that will occur in the environment: color type
+color is the color of the box. Can be “red”, “green”, “blue”, “purple”, “yellow” or “grey”.
+type is the type of the object. Can be “ball”, “box” or “key”.
+There will be one and only one red ball. The red ball may be present anywhere in the environment, and it may or may not be visible in a given frame.
+The agent appears as a red triangle positioned at the center of the bottom row, always facing upward.
+The edge of the visible world is flanked by a uniform grey strip, which separates a lighter interior grid from a darker exterior region.
+No objects appear beyond this strip.
+There are no doors in this environment.
+The observation is ego-centric: the agent is always placed at the center of the bottom row in the image.
+It can see all tiles directly in front of it, as well as up to three tiles to each side—limited only by the surrounding walls.
+The image consistently covers a 7×7 grid area, centered horizontally around the agent’s position.
+You need to control the agent to navigate to the red ball.
+The agent cannot move onto a grid cell that contains an object or a wall.
+As a result, any attempt to enter an occupied cell or pass through a wall is invalid and should be avoided.
+Any attempt to turn towards an occupied cell should also be avoided unless the agent has long-term intentions (e.g.
+making an U-turn to go back).
+BabyAI uses an action space of Discrete(7), but only the first three actions are used:
+Value | Meaning
+   0   | Turn left
+   1   | Turn right
+   2   | Move forward
+   3-6 | Unused
+A.5 Appendix E: VLM Hallucination Analysis
+Intuitively, one might also consider using VLMs to label features for a dataset of images.
+While VLMs may not provide precise positional values, coarse ordinal spatial descriptors, such as "left," "far left," or "right", can be sufficient for environments like skiing.
+However, our experiments demonstrate that this approach is fundamentally limited.
+Most VLMs visual backbones, such as CLIP or SigLIP, encode images by dividing them into fixed-size patches (e.g., 32×32 pixels).
+Each patch is linearly embedded and processed through transformer layers to model contextual relationships across patches.
+However, this patch-based architecture struggles to maintain accurate spatial relationships between objects when multiple objects fall within the same patch.
+This is a critical failure for feature extraction, as relevant features can appear anywhere in the image, and distinguishing between multiple objects within a single patch becomes impossible, even when grid markers are added to disambiguate spatial zones.
+To illustrate this issue, we conducted a simple experiment (Figure 10). When prompted with “What do you see? Describe tile by tile,” chatgpt-4o-latest responded:
+  • First tree (smaller): around x2y15 (bottom-left)
+  • Second tree (larger): around x4y13 (just above the first tree)
+However, the actual spatial relationship between grid x2y15 and x4y13 is not consistent with the relative positioning described by the model.
+Figure 10: Hallucination analysis on labeled grids.

### v3

@@
 Abstract
-Semantic interpretability in Reinforcement Learning (RL) enables transparency and verifiability by making the agent’s decisions understandable and verifiable.
-Achieving this, however, requires a feature space composed of human-understandable concepts, which traditionally rely on human specification and may fail to generalize to unseen environments.
-We introduce interpretable Tree-based Reinforcement learning via Automated Concept Extraction (iTRACE), an automated framework that leverages pre-trained vision-language models (VLM) for semantic feature extraction and interpretable tree-based models for policy optimization.
-iTRACE first extracts semantically meaningful features, then maps them to policies via interpretable trees.
+Semantic interpretability in Reinforcement Learning (RL) enables transparency and verifiability of decision-making.
+Achieving semantic interpretability in reinforcement learning requires (1) a feature space composed of human-understandable concepts and (2) a policy that is interpretable and verifiable.
+However, constructing such a feature space has traditionally relied on manual human specification, which often fails to generalize to unseen environments.
+Moreover, even when interpretable features are available, most reinforcement learning algorithms employ black-box models as policies, thereby hindering transparency.
+We introduce interpretable Tree-based Reinforcement learning via Automated Concept Extraction (iTRACE), an automated framework that leverages pre-trained vision-language models (VLM) for semantic feature extraction and train a interpretable tree-based model via RL.
 To address the impracticality of running VLMs in RL loops, we distill their outputs into a lightweight model.
-By leveraging Vision-Language Models (VLMs) to automate tree-based reinforcement learning, iTRACE eliminates the need for human annotation traditionally required by interpretable models, while also addressing the limitations of VLMs alone, such as their lack of grounding in action spaces and inability to directly optimize policies.
-iTRACE outperforms MLP baselines that use the same interpretable features and matches the performance of CNN-based policies, producing verifiable, semantically interpretable, and human-aligned behaviors without requiring human annotation.
-1 Introduction
+By leveraging Vision-Language Models (VLMs) to automate tree-based reinforcement learning, iTRACE loosens the reliance the need for human annotation that is traditionally required by interpretable models.
+In addition, it addresses key limitations of VLMs alone, such as their lack of grounding in action spaces and their inability to directly optimize policies.
+We evaluate iTRACE across three domains: Atari games, grid-world navigation, and driving.
+The results show that iTRACE outperforms other interpretable policy baselines and matches the performance of black-box policies on the same interpretable feature space.
+I Introduction
 Reinforcement learning (RL) agents operate in dynamic environments, often developing policies that rely on latent representations not directly interpretable by humans [1].
-Ensuring that these learned representations align with human-understandable concepts, a goal often referred to as semantic interpretability, is crucial for ensuring safety and preventing unintended harmful actions.
-While progress has been made in developing interpretable AI models, achieving semantic interpretability in RL remains particularly challenging  [22].
-Semantic interpretability refers to the extent to which a model’s features, decisions, or internal representations can be understood and explained in terms of human-understandable concepts.
-Achieving this requires two essential components: (1) the inputs to the decision-making model must be intelligible to facilitate downstream interpretation, and (2) the decision-making process must be built with an interpretable model that allows humans to understand and explain its rationale  [9].
-In essence, achieving semantic interpretability involves mapping high-dimensional input data to human-understandable features that can serve as a low-dimensional input space for interpretable decision-making models.
-To learn an interpretable model on a semantically interpretable feature space, various model classes can be considered, including decision trees and rule-based policies, both of which facilitate explicit reasoning about an agent’s actions.
-Decision trees have been used to represent policies in a structured, human-readable format [30, 26, 18].
+Ensuring the learned representations align with human-understandable concepts, a goal often referred to as semantic interpretability, is crucial for ensuring safety and preventing unintended harmful actions.
+While progress has been made in interpretable AI, achieving semantic interpretability in RL remains particularly challenging [2].
+It requires (1) constructing a feature space composed of human-understandable concepts, a process that has traditionally relied on human annotation and lacks generalizability.
+(2) In addition, it requires learning policies that are both high-performing and transparent [3].
+To learn an interpretable model on a semantically interpretable feature space, various model classes can be considered, including decision trees and rule-based policies, both of which facilitate explicit reasoning about an agent’s actions [3].
+Decision trees have been used to represent policies in a structured, human-readable format [4, 5].
 Rule-based methods offer an alternative by encoding policies in a form that can be explicitly followed and verified [6, 7].
 However, none of these prior works address high-dimensional image or video inputs; instead, they rely either on human-defined semantic concepts or ground-truth feature vectors provided by the environment.
-Existing approaches have attempted to bridge the gap between raw image inputs and interpretable features [2] to facilitate model understanding but fall short in semantic interpretability.
-Many approaches either produce mathematical or visual representations that do not correspond to human-understandable concepts [15, 10] or rely on predefined feature spaces that aren’t generalize across different environments, using CNNs [8] or ViTs [23] to establish mappings between inputs and semantic features.
-As a result, human intervention has traditionally played a key role, either by manually defining semantically meaningful feature spaces [40] or by labeling machine-generated features to associate them with semantically meaningful concepts [29].
-The emergence of Vision-Language Models (VLMs) has opened new possibilities for overcoming these challenges.
-These models, known for their capabilities in fine-grained visual understanding, language reasoning, and decision-making  [4, 38, 6], have seen widespread use in various tasks  [38, 39].
-In this work, we demonstrate that VLMs can substitute humans in defining and extracting semantic features, serving not merely as a cost-effective fallback to human labor, but as a scalable and automated solution that matches human performance.
-We propose interpretable Tree-based Reinforcement learning via Automated Concept Extraction (iTRACE), which is to the best of our knowledge the first automated framework that achieves semantic interpretability in RL and generalizes effectively to unseen environments.
-By leveraging VLMs to extract features from raw environment inputs and training with Interpretable Control Trees, we manage to provide the rationale behind a RL algorithm in terms of human-understandable concepts.
-Notably, our framework (i) maintains competitive performance with black-box models while surpassing uninterpretable models built on the same feature space, (ii) supports both image and video inputs, effectively capturing temporal information, (iii) generalizes to unseen environments without human-in-the-loop, and (iv) enables computational efficiency by training a lightweight vision model as an alternative to VLM-based feature extraction.
-2 Related work
-2.1 Interpretable reinforcement learning
+Figure 1: High-level overview of iTRACE.
+In this paper, we tackle the challenge of mapping high-dimensional observations into human-understandable features that can serve as a low-dimensional input space for interpretable decision-making models.
+Prior work has attempted to bridge the gap between raw image inputs and interpretable features [8], but existing methods fall short of true semantic interpretability.
+Some approaches generate mathematical or visual representations that do not correspond to human-understandable concepts [9, 10].
+Others rely on predefined feature spaces that fail to generalize across environments [11, 12].
+As a result, human involvement remains central: researchers either manually define semantic features [13] or label machine-generated ones [14].
+However, this reliance on human annotation is not automated and incurs significant time costs, limiting scalability to new domains.
+The emergence of Vision-Language Models (VLMs) has opened new possibilities for addressing these challenges, particularly the difficulty of building human-aligned feature extractor.
+VLMs, known for their capabilities in visual understanding, language reasoning, and decision-making  [15], have seen widely used in various tasks  [15, 16].
+In this work, we propose interpretable Tree-based Reinforcement learning via Automated Concept Extraction (iTRACE).
+iTRACE leverages VLMs to extract semantic features from raw pixel inputs and trains Interpretable Control Trees to provide the rationale behind RL decisions in terms of human-understandable concepts.
+Our framework (i) achieves performance competitive with black-box models while surpassing uninterpretable models built on the same feature space, (ii) supports both single-frame and multi-frame inputs, enabling the capture of temporal information, (iii) generalizes to unseen environments without human involvement, and (iv) ensures computational efficiency by training a lightweight vision model as an alternative to VLM-based feature extraction.
+To the best of our knowledge, iTRACE is the first automated framework to achieve semantic interpretability in RL while generalizing effectively to new environments without human-in-the-loop.
+II Related work
+II-A Interpretable reinforcement learning
 Figure 2: Overview of iTRACE, our proposed framework.
 A.
-Feature Extraction involves a three-step process where: (1) The VLM identifies key features given domain knowledge and environment observation.
-(2) The VLM generates code to extract features from a segmentation of a single image frame and determines whether certain key features require multiple frames.
-If so, it extends the single-frame extraction code to handle multi-frame inputs.
-The segmentation is performed using the Segment Anything Model (SAM).
-(3) An exploration policy is used to generate a dataset of feature–raw image frame pairs by applying SAM and the feature extraction code.
-To avoid the computational overhead of running SAM at every reinforcement learning step, we train an Extractor Module as an efficient substitute.
+Feature Extraction follows three steps: (1) the VLM identifies key semantic features from domain knowledge and observations; (2) it generates code to extract features from SAM-based segmentations, extending to multi-frame inputs if needed; and (3) an exploration policy generates feature–image pairs, which are used to train a lightweight Extractor Module to replace costly SAM queries during RL.
 B.
-Tree-based RL Training leverages the pretrained Extractor Module to obtain interpretable features from image observations, which are subsequently used as inputs to the policy network (an Interpretable Control Tree) to generate control actions for the RL agent.
-Interpretable RL is inherently challenging compared to standard supervised learning, as it needs to account for the long-term consequences of actions, partial observability, and sparse feedback during exploration [22].
-Post-hoc methods like Jacobian-based, perturbation-based, and attribution-based saliency maps  [25, 10, 34] offer limited interpretability, as they lack transparency in both inputs and decision-making [9].
-Approaches to achieve intelligible decision-making include rule lists and decision trees [9], among which decision trees are arguably the easiest to verify  [27].
-Among state-of-the-art tree methods, Differentiable Decision Trees (DDTs) support gradient-based optimization of interpretable models [27], while Interpretable Continuous Control Trees (ICCTs) extend this capability to continuous control tasks, balancing performance with interpretability [19].
-Similarly, POETREE (Policy Extraction through Decision Trees) [17] introduces dynamically growing probabilistic decision trees that adapt throughout training, allowing the tree complexity to evolve based on the task requirements.
-Our work addresses the limitation of prior tree-based methods, which previously required manual feature engineering or were limited to structured inputs
-2.2 Feature extraction for reinforcement learning
+Tree-based RL Training then leverages the Extractor Module to obtain interpretable features, which serve as inputs to an Interpretable Control Tree for policy learning.
+Interpretable RL remains an open challenge, as current methods struggle to combine human-understandable policies with high performance, to determine when such policies are feasible, and to maintain transparency in high-dimensional state spaces [2].
+Following Rudin’s definition, we view Interpretable RL as learning constrained models whose reasoning humans can directly understand, distinct from post-hoc explainability tools such as SHAP [17], LIME [18], or saliency maps [19, 20], which are often misleading and susceptible to hallucinations.
+Our work addresses the limitation of prior tree-based methods, which previously required manual feature engineering or were limited to structured inputs.
+Among inherently interpretable models, tree-based methods stand out for producing explicit human-readable rules [21], handling categorical and nonlinear interactions [22], supporting sparsity constraints [23], and achieving accuracy competitive with black-box models when properly optimized [24, 25, 26].
+Recent advances extend these strengths: Differentiable Decision Trees (DDTs) enable gradient-based optimization of interpretable models [27]; Interpretable Continuous Control Trees (ICCTs) extend this approach to continuous control tasks, balancing transparency and performance [28]; and POETREE introduces dynamically growing probabilistic trees that evolve in complexity during training [29].
+II-B Feature extraction for reinforcement learning
 Nonparametric methods (e.g., manifold and spectral learning) and nonlinear parametric methods (e.g., deep learning) are commonly used in RL for feature extraction but typically do not produce interpretable features [9].
-Other work focuses on the extraction of visual features that improve performance but do not guarantee human-interpretable concepts.
-Yet, few works attempt to construct semantically meaningful feature spaces without human supervision for RL training.
-[35] leverages information from Atari instruction manuals to construct the feature space, then uses unsupervised object detection and grounding to acquire the corresponding objects in the environment.
-However, they report that the extracted features could be noisy and inconsistent  [35].
-Our framework differs from prior work by enabling automatic feature definition and extraction without human-in-the-loop.
-2.3 VLMs for reinforcement learning
-VLMs have been applied in various aspects of RL training, including reward design, preference feedback, and representation extraction [21, 33, 6].
-One prominent approach involves using VLMs to generate reward functions.
-[21] use VLMs as zero-shot reward models to specify tasks via natural language, while  [32] leverages VLMs to produce dense reward functions through code generation.
-Additionally,  [33, 31] use VLMs to provide preference feedback with which they updated the reward functions for preference-based RL.
-Another line of work leverages VLMs as feature extractors to obtain valuable information which facilitates training.
-[6] extract promptable representations from VLMs and decode output tokens into embeddings, which are used to initialize RL policies.
-However, these representations are not guaranteed to be accurate or semantically meaningful.
-To our knowledge, this is the first approach that uses VLMs to automatically construct semantically interpretable feature spaces for RL.
-3 Methodology
-We introduce interpretable Tree-based Reinforcement learning via Automated Concept Extraction (iTRACE), a framework designed to automatically learn an interpretable decision-making RL policy from an unseen environment.
-iTRACE leverages the visual comprehension and reasoning capabilities of VLMs automate feature extraction through a three-step pipeline: (1) it employs zero-shot prompting on the environment to identify a set of semantic features that deterministically contribute to the decision-making, (2) it generates feature extraction codes that captures both static and temporal information, and (3) it collects a dataset with VLM-generated extraction codes to train a lightweight vision model (which we refer to as the Extractor Module) as a computationally efficient substitute.
-Subsequently, utilizing the Extractor Module to replicate VLM feature extraction, iTRACE performs tree-based RL training with an Interpretable Control Tree, establishing a transparent mapping between the semantic feature space and action outputs to derive a semantically interpretable policy for decision-making.
-3.1 Feature extraction
-The feature extraction process of iTRACE involves a three-step pipeline, as shown in Fig.
-1.
-In Step 1, we query the VLM to generate key features that contribute to the decision in the environment.
-In Step 2, we query the VLM to generate code that extracts the identified features, either from a single image or a list of images depending on whether the features demand temporal context.
-Both raw observations and segmentation masks are provided to the VLM in this process.
-In Step 3, we collect a dataset from the environment with the VLM-generated extractor code, with which we train an Extractor Module to increase program speed in RL training.
-The methodology is formalized below.
-Step 1: Querying the VLM for Semantic Features.
-To extract intelligible features from an unseen environment that could be used to train an effective RL policy, we need to first determine what features are useful to model decision-making.
+Other approaches focuses on identifying visual representations that boost performance without guaranteeing human-meaningful concepts: SSINet highlights task-relevant regions with attention masks [19], while perturbation-based saliency maps identify influential pixels [10].
+Yet, few works construct semantically meaningful feature spaces without human supervision for RL training.
+For example, Wu et al.
+[30] automatically parse features from Atari manuals, but their extracted features remain noisy and still rely on oracle human-defined information.
+In contrast, our framework automatically defines and extracts features without human-in-the-loop assumptions.
+III Interpretable Feature Extraction
+In this section, we introduce the interpretable feature extraction component of the iTRACE framework, shown in the first block of Fig.
+2.
+This component leverages the visual comprehension and reasoning capabilities of VLMs to automate feature extraction through a three-step pipeline: (1) zero-shot prompting is used to identify semantic features that deterministically influence decision-making (line 3 in Algorithm 1); (2) feature extraction code is generated to capture both static and temporal information (line 4 in Algorithm 1).
+Both raw observations and segmentation masks are provided to the VLM in this process; and (3) a dataset is collected using the VLM-generated code to train a lightweight vision model (the Extractor Module) as a computationally efficient substitute (line 5 and line 6 in Algorithm 1).
+Step 3 is crucial because VLM-generated code depends on image segmentation, which is prohibitively slow and non-batchable when performed at every RL step, resulting in impractical training times.
+III-A Deduce features
+To extract intelligible features from an unseen environment that could be used to train an effective RL policy, we need to first determine what features are useful to model decision-making (line 3 in Algorithm 2).
 In this step, we provide the VLM with:
-  • $\mathcal{D}$: Basic domain knowledge about the environment.
+  • $\mathcal{D}$: Domain knowledge limited to basic information, obtainable via foundation model queries or agentic exploration (Sec. VI).
   • $\mathcal{I}_{0}$: A diverse list of example image observations from the environment.
@@
 $\mathcal{F}=\mathcal{F}_{s}\cup\mathcal{F}_{m},\quad\mathcal{F}=\mathcal{M}(\mathcal{D},\mathcal{I}_{0}).$ (1)
-For example, in the Skiing environment, $\mathcal{F}_{s}$ includes the positions of the next gate (left pole x, right pole x, and y), the closest tree (x, y) and the skier’s x-position, while $\mathcal{F}_{m}$ includes the horizontal speed of the skier.
-We instruct the VLM to provide a "requires_temporal_context" field for each feature, enabling it to automatically distinguish between static and temporal features and assign them to $\mathcal{F}_{s}$ or $\mathcal{F}_{m}$.
+For example, in the Skiing environment, as deduced by the VLM, $\mathcal{F}_{s}$ includes the positions of the next gate (left pole x, right pole x, and y), the closest tree (x, y) and the skier’s x-position, while $\mathcal{F}_{m}$ includes the horizontal speed of the skier.
+We instruct the VLM to give a "temporal_context" boolean field for each feature, enabling it to automatically distinguish between static and temporal features and assign them to $\mathcal{F}_{s}$ or $\mathcal{F}_{m}$.
 While $\mathcal{F}_{m}$ may be unnecessary in some environments, it is essential in others.
@@
 We allow the VLM to determine whether temporal context is required for effective decision-making in a given environment.
-A complete list of extracted features is provided in Appendix A.2.2.
-Step 2: Querying the VLM for Feature Extraction Code.
+III-B Generate feature extraction code
 In this step, we query the VLM to generate feature extraction code, denoted by $\mathcal{G}$, to retrieve the identified set of features $\mathcal{F}$ from one or multiple image-segmentation pairs $(\mathcal{I},\mathcal{S})$.
-Recognizing that hallucination is a well-documented challenge in generative VLM models [16] (see Appendix.
-A.5), we use segmentation to provide accurate spatial information and ground the generated extraction code in accurate visual data.
+Recognizing that hallucination is a well-documented challenge in generative VLM models [31], we use segmentation to provide accurate spatial information and ground the generated extraction code in visual data.
 This strategy aligns with common practices in the field [32, 33].
-The visual segmentation $\mathcal{S}$ is computed using Segment Anything [13], denoted by $Seg$, such that $\mathcal{S}=\text{Seg}(\mathcal{I})$.
+The visual segmentation $\mathcal{S}$ is computed using Segment Anything Model (SAM) [34], denoted by $Seg$, such that $\mathcal{S}=\text{Seg}(\mathcal{I})$.
 We use the ViT-H SAM model with its default parameters and no modifications to ensure the generalizability of our approach.
 Example Segmentations Labeling: While segmentation improves spatial grounding, its interpretation by the VLM remains prone to hallucination.
-To ensure the generated code generalizes across diverse environment frames, the VLM must be exposed to a representative set of example images.
-However, the raw JSON format of segmentation can lead to confusion between segments across different images.
+To ensure the generated code generalizes across diverse environments, the VLM must be exposed to a representative set of example images.
+However, providing multiple segmentation JSONs at once can cause the VLM to conflate objects across images, consistent with known failures of VLMs to maintain object identity across contexts [35].
 To address this, we first prompt the VLM to label a small set of segmentations associated with example frames, processing one image at a time.
 An example of this labeling process is shown in Fig.
-9(b), where all segment labels are generated by the VLM based on raw segmentation masks.
+4(b), where all segment labels are generated by the VLM based on raw segmentation masks.
 Following this, depending on whether the feature set requires temporal context, the VLM is instructed to take one or multiple image-segmentation pairs $(\mathcal{I},\mathcal{S})$.
 The labeled segmentations serve as reference examples, helping the VLM infer the most likely attribute values of relevant objects and formulate logic to filter out noisy segments.
-The code generation process involves two steps:
-Single Image Extraction: For features in $\mathcal{F}_{s}$, the VLM generates code $\mathcal{G}_{s}$ to extract their values from a single image-segmentation pair $(\mathcal{I},\mathcal{S})$.
-To enhance the robustness of the generated code, the VLM is provided with multiple example images and their labeled segmentation during this stage to ensure it has been exposed to most possible scenarios.
-See Algorithm 1.
-Also see Fig.
-9(c) for an example output on segmentation masks in the Skiing environment.
-Multiple Image Extraction: For features in $\mathcal{F}_{m}$, if any, the VLM generates code $\mathcal{G}_{m}$, which extends $\mathcal{G}_{s}$ to process sequential image-segmentation pairs $\{(I_{t-i},S_{t-i}),\dots,(I_{t},S_{t})\}$.
-Multiple image-segmentation pairs $(\mathcal{I},\mathcal{S})$ are provided to the VLM during generation.
-See Algorithm 2.
-The final feature extraction code $\mathcal{G}$ is defined in Eq. 2:
-$\mathcal{G}=\begin{cases}\mathcal{G}_{s},&\text{if }\mathcal{F}_{m}=\emptyset;\\ \mathcal{G}_{m}(\mathcal{G}_{s}),&\text{if }\mathcal{F}_{m}\neq\emptyset.\end{cases}$ (2)
-Algorithm 1 Generating Extraction Code: Single Image
-1:  Input: Feature set $\mathcal{F}$, images $\{\mathcal{I}_{i}\}_{i=1}^{n}$, VLM $\mathcal{M}$
-2:  Initialize $\mathcal{G}_{s}=null$
-3:  $examples\leftarrow\emptyset$
-4:  for each image $\mathcal{I}_{i}$ in $\{\mathcal{I}_{i}\}_{i=1}^{n}$ do
-5:     $\mathcal{S}_{i}\leftarrow\text{Seg}(\mathcal{I}_{i})$
-6:     Add $(\mathcal{I}_{i},\mathcal{S}_{i})$ to $examples$
-7:  end for
-8:  $\mathcal{G}_{s}\leftarrow\mathcal{M}_{\text{Generate}}(\mathcal{F},examples)$
-Algorithm 2 Generating Extraction Code: Multiple Images
-1:  Input: Feature set $\mathcal{F}$, images $\{\mathcal{I}_{i}\}_{i=1}^{n}$, $\mathcal{G}_{s}$, VLM $\mathcal{M}$
-2:  Initialize $\mathcal{G}_{m}=null$
-3:  $examples\leftarrow\emptyset$
-4:  for each image $\mathcal{I}_{i}$ in $\{\mathcal{I}_{i}\}_{i=1}^{n}$ do
-5:     $\mathcal{S}_{i}\leftarrow\text{Seg}(\mathcal{I}_{i})$
-6:     Add $(\mathcal{I}_{i},\mathcal{S}_{i})$ to $examples$
-7:  end for
-8:  $\mathcal{F}_{m}\leftarrow{\mathcal{M}_{\text{JudgeMultiFrame}}}(\mathcal{F})$
-9:  if $\mathcal{F}_{m}\neq\emptyset$ then
-10:     $\mathcal{G}_{m}\leftarrow\mathcal{M}_{\text{Generate}}(\mathcal{F},examples,\mathcal{G}_{s})$
-11:  end if
-12:  Return: $\mathcal{G}_{m}$
+The code generation process is described in line 4-9 in Algorithm 2), where the VLM first generates code to extract values from a single image–segmentation pair $(\mathcal{I},\mathcal{S})$, and then generates code $\mathcal{G}_{m}$, extending $\mathcal{G}_{s}$ to handle sequences of image-segmentation pair $\{(I_{t-i},S_{t-i}),\dots,(I_{t},S_{t})\}$.
+Algorithm 1 iTRACE Overall Framework
+1:Input: Environment $\mathcal{E}$, VLM $\mathcal{M}$, RL algorithm (PPO), horizon $H$
+2: // Call perception stack
+[…]
+[…]
+[…]
+[…]
+7: // Call RL stack
+[…]
+9:Return: trained ICT policy $\pi_{\theta}$
+Algorithm 2 Perception Stack: Automated Feature Extraction
+1:Input: Environment $\mathcal{E}$, VLM $\mathcal{M}$, Domain $\mathcal{D}$
+2: Sample example images $\{I_{i}\}_{i=1}^{n}$ from $\mathcal{E}$ // Step 1: Deduce features
+[…]
+[…]
+5:if $F_{m}\neq\emptyset$ then
+6:  $G_{m}\leftarrow\mathcal{M}_{\text{Generate}}(F_{m},G_{s})$
+7:  $G\leftarrow(G_{s},G_{m})$
+8:else
+9:  $G\leftarrow G_{s}$
+10:end if// Step 3: Train extractor module
+11: Collect dataset $\mathcal{T}=\{(I_{t},G(I_{t}))\}$ via exploration
+12: Train lightweight extractor $\phi$ on $\mathcal{T}$
+13:Return: Extractor module $\phi$
+Algorithm 3 RL Stack: Training Interpretable Control Tree
+1:Input: Extractor module $\phi$, environment $\mathcal{E}$, horizon $H$, PPO parameters
+2: Initialize ICT policy $\pi_{\theta}$
+3:for episode $1,\dots,N$ do
+4:  Reset $\mathcal{E}$, observe raw frame $I_{0}$
+5:  for t = 1 to H do
+6:   $f_{t}\leftarrow\phi(I_{t})$ // semantic features
+7:   $a_{t}\sim\pi_{\theta}(f_{t})$
+8:   Apply $a_{t}$ in $\mathcal{E}$, receive $I_{t+1},r_{t}$
+9:  end for
+10:  Update $\pi_{\theta}$ via PPO on collected $(f_{t},a_{t},r_{t})$
+11:end for
+12:Return: trained ICT policy $\pi_{\theta}$
 (a) Original image observation.
@@
 Figure 3: An example of how raw image observations are processed in our pipeline.
-(a) Raw image observations are segmented using Segment Anything.
+(a) Raw image observations are segmented using SAM.
 (b) A VLM assigns semantic labels to segments in a small set of example images.
@@
 Step (b) is only performed on the example set to guide code generation.
-Step 3: Dataset Generation and Training the Extractor Module.
-In this step, we execute $\mathcal{G}$ on trajectories collected from the environment to generate the feature dataset.
-For each trajectory, a series of images $\mathcal{T}=\{I_{1},I_{2},\dots,I_{k}\}$ is generated, with corresponding visual segmentations $\mathcal{S}_{t}=\text{Seg}(I_{t})$ for each $t$.
-The extractor code $\mathcal{G}$ processes each image-segmentation pair $(I_{t},S_{t})$ independently to extract features $\mathbf{f}_{t}=\{f_{1},f_{2},\dots,f_{n}\}$.
-The dataset, consisting of image-feature pairs $\{(\mathcal{I}_{t},\mathbf{f}_{t})\}$, is then used to train an Extractor Module as a batch-operable and efficient replacement to $\mathcal{G}_{s}$.
+III-C Dataset generation and training the extractor module.
+In this step, we address the inefficiencies of the VLM-generated feature extraction code for single-image frames, $\mathcal{G}_{s}$, by training an lightweight Extractor Module as its replacement.
+While $\mathcal{G}_{s}$ can be executed directly on the dataset, the processing pipeline is inherently non-batchable, as it handles image segmentations independently.
+This non-batchability severely impacts downstream RL training.
+We execute $\mathcal{G}$ on trajectories collected from the environment to generate the feature dataset.
+For each trajectory, a series of images, $\mathcal{T}=\{I_{1},I_{2},\dots,I_{k}\}$ is generated, with corresponding visual segmentations $\mathcal{S}_{t}=\text{Seg}(I_{t})$ for each $t$.
+The extractor code $\mathcal{G}$ processes each image-segmentation pair $(I_{t},S_{t})$ independently to extract features $\mathbf{f}_{t}=\{f_{1},f_{2},\dots,f_{n}\}$ (line 11 in Algorithm 2).
+The dataset, consisting of image-feature pairs $\{(\mathcal{I}_{t},\mathbf{f}_{t})\}$, is then used to train an Extractor Module as a batch-operable and efficient replacement to $\mathcal{G}_{s}$ (line 12 in Algorithm 2).
 The trained model predicts semantic features directly from raw images, enabling significantly faster feature extraction during RL training.
-In our experiments, we use a CNN-based Extractor Module (architecture included in Appendix A.3) due to its low computational overhead in simulated environments.
-However, the framework is compatible with any vision model, including fine-tuned ViT transformers or other task-specific architectures based on deployment needs.
-If some features require temporal context (i.e., $\mathcal{F}_{m}\neq\emptyset$), $\mathcal{G}_{m}$, consisting of simple mathematical operations on the outputs of $\mathcal{G}_{s}$, is manually adapted to utilize the Extractor Module in place of $\mathcal{G}_{s}$.
-For example, in the Skiing environment, $\mathcal{G}_{m}$ calculates skier horizontal speed by subtracting the positions returned by $\mathcal{G}_{s}$ The manual adaptation preserves the exact same logic while ensuring that the subtraction operation is properly batched, significantly accelerating the training process by several orders of magnitude.
-3.2 Tree-based RL training
-After obtaining the Extractor Module, we train an interpretable tree-based model with RL to connect the semantic features with action outputs.
-During RL training, our framework utilizes the pre-trained Extractor Module to extract semantic features from image observations, as shown in Fig.
-1.
+In our experiments, we use a CNN-based Extractor Module due to its low computational overhead in simulated environments.
+While the framework is compatible with any lightweight vision model, including fine-tuned ViTs or task-specific architectures, CNNs provide an effective and widely validated baseline that balances accuracy, efficiency, and robustness.
+The execution time comparison between VLM-generated feature extraction code and our CNN-based Extractor Module is demonstrated in Table I.
+TABLE I: Runtime speedup on Skiing. Similar results achieved on other domains.
+Batch VLM Code CNN Speedup
+[…]
+[…]
+[…]
+[…]
+[…]
+If some features require temporal context (i.e., $\mathcal{F}_{m}\neq\emptyset$), $\mathcal{G}_{m}$, consisting of simple mathematical operations on the outputs of $\mathcal{G}_{s}$, is adapted to utilize the Extractor Module in place of $\mathcal{G}_{s}$.
+For example, in the Skiing environment, $\mathcal{G}_{m}$ calculates skier horizontal speed by subtracting the positions returned by $\mathcal{G}_{s}$
+IV Tree-based RL training
+Leveraging the semantic features extracted from the raw visual inputs courtesy of the perception stack, we now focus on training an interpretable RL model that maps semantic features to action outputs, as shown in Fig.
+2.
 The policy network takes these semantic features as input and predicts a distribution over actions, from which an action is sampled and applied to the controlled agent at each time step.
-Unlike traditional RL algorithms that use black-box models as the policy network, we use an Interpretable Control Tree (ICT) as our policy network to achieve transparent decision-making and human-aligned reasoning while remaining differentiable for end-to-end optimization.
-3.2.1 Interpretable control tree
-Interpretable Control Tree (ICT) is adapted from Interpretable Continuous Control Tree (ICCT) proposed by Paleja et al.
-[18].
-The ICT is structured similarly as ICCT, where each decision node determines the routing of input features, $\mathbf{x}$, to sparse linear controllers at the leaf nodes.
+Unlike traditional RL algorithms that use black-box models as the policy network, we use an Interpretable Control Tree (ICT) adapted from Interpretable Continuous Control Tree (ICCT) proposed by Paleja et al.
+[5] as our policy network to achieve transparent decision-making and human-aligned reasoning while remaining differentiable for end-to-end optimization.
+Unlike ICCT, ICT supports both discrete and continuous actions without re-architecture, which helps generalizability.
+In the discrete case, leaf nodes in ICT output the possibility of taking each action as a Probability Mass Function (PMF).
+If there are six possible actions in the discrete action space, then there will be six sparse linear functions of the input features followed by a softmax function at each leaf node.
+And for the continuous action space, leaf nodes output the Probability Density Function (PDF) over possible actions.
+The detailed RL training is listed in Algorithm 3
+IV-A Interpretable control tree
+Each decision node in ICT determines the routing of input features, $\mathbf{x}$, to sparse linear controllers at the leaf nodes.
 Each decision node, $i$, is parameterized by weights $\mathbf{w}_{i}$, a bias $b_{i}$, and a steepness parameter $\alpha$ to output a fuzzy Boolean in $[0,1]$, as defined in Eq. 2.
@@
 $k=\arg\max_{j}\left|w_{i}^{j}\right|$ (3)
-Since the $\arg\max$ operator is non-differentiable, we approximate this function with the Gumbel-Softmax  [11], as per Eq.
-5, where $g_{j}\sim\text{Gumbel}(0,1)$ and $\tau$ is a temperature parameter.
+Since the $\arg\max$ operator is non-differentiable, we approximate this function with the Gumbel-Softmax  [36], as per Eq. 4, where $g_{j}\sim\text{Gumbel}(0,1)$ and $\tau$ is the temperature.
 $f(w_{i})^{k}=\frac{\exp\left(\frac{w_{i}^{k}+g_{k}}{\tau}\right)}{\sum_{j}\exp\left(\frac{w_{i}^{j}+g_{j}}{\tau}\right)}$ (4)
-During inference, a hard threshold is applied to enforce discrete decision-making using only one feature per node, as per Eq. 6.
+During inference, a hard threshold is applied to enforce discrete decision-making using only one feature per node, as per Eq. 5, where $\mathbf{1}$ is the indicator function.
 $y_{i}=\mathbf{1}(\alpha(w_{i}^{k}x^{k}-b_{i})>0)$ (5)
-While during training, gradients are allowed to flow through discrete decisions using the straight-through trick [3].
-This approach ensures that each decision node remains interpretable while allowing for end-to-end gradient-based learning.
+During training, gradients flow through discrete decisions using the straight-through trick [37].
+This approach ensures each decision node remains interpretable while allowing end-to-end gradient-based learning.
 Each leaf node, $\ell_{d}$, in the ICT is parameterized by a set of sparse linear controllers, as formulated in Eq.
-7, where $\mathbf{\beta}_{d}$ is a sparsely activated weight vector, $\phi_{d}$ is a bias term, and $\mathbf{u}_{d}$ is a $k$-hot encoding vector that enforces sparsity, computed in Eq.
-8.
+6, where $\mathbf{\beta}_{d}$ is a sparsely activated weight vector, $\phi_{d}$ is a bias term, and $\mathbf{u}_{d}$ is a $k$-hot encoding vector that enforces sparsity, computed in Eq.
+7.
 The per-leaf selector weight, $\theta_{d}$, learns the relative importance of different input features.
@@
 $\mathbf{u}_{d}=h\left(\mathrm{softmax}\left(|\mathbf{\theta}_{d}|\right)\right),$ (7)
-Unlike ICCT, ICT supports both discrete and continuous actions without re-architecture, which helps generalizability.
-In the discrete case, leaf nodes in ICT output the possibility of taking each action as a Probability Mass Function (PMF).
-If there are six possible actions in the discrete action space, then there will be six sparse linear functions of the input features followed by a softmax function at each leaf node.
-And for the continuous action space, leaf nodes output the Probability Density Function (PDF) over possible actions.
-4 Experiments
-4.1 Environment
-We evaluate our approach across three diverse environments: Atari Skiing, a fast-paced pixel-art game requiring gate navigation; BabyAI-GoToRedBall, a symbolic grid world for instruction following; and OpenAI Gym Highway, a 2D environment for high-level autonomous driving decisions like lane switching and speed control.
+V Experiments
+V-A Environment
+We evaluate iTRACE in three diverse environments chosen to ensure the VLM performs genuine reasoning rather than relying on memorized knowledge [38]: Atari Skiing, a pixel-art game; BabyAI-GoToRedBall, a symbolic grid world for navigation; and OpenAI Gym Highway, a 2D driving environment.
+While BabyAI and Highway provide default feature vectors, these describe the full environment state rather than the task-relevant semantic features required for decision making.
 The Atari Skiing environment provides raw image state spaces without any pre-defined feature vectors or feature sets.
 BabyAI and Highway offer default feature vectors, but these vectors describe the full state of the environment rather than focusing on the task-relevant semantic features needed for decision making.
-By selecting these environments, we ensure that the VLM engages in genuine reasoning over the environment to identify key semantic features, rather than relying on prior internet-based knowledge, a phenomenon referred to as training data memorization [5].
-4.2 Feature extraction
-4.2.1 VLM deducing key semantic features.
+V-B Feature extraction
+V-B1 VLM deducing key semantic features.
 For this step, we utilize ChatGPT-4o-latest, one of the current leading models in language reasoning.
@@
 For all three environments, we query the VLM 50 times and use majority vote to get the most prominent answer from the VLM.
-A complete set of those features are included in Appendix A.2.2.
-4.2.2 VLM generating Feature Extraction code.
-For this step, we utilize claude-3.5-sonnet-20241022, the current leading model in code generation.
+A complete set of those features are included in Table II.
+TABLE II: Feature schemas for BabyAI RedBall, Highway, and Skiing environments. Type: C = Continuous, B = Binary. Temporal: T = True, F = False.
+Feature Type Temp.
+BabyAI
+red_ball.{x,y}_position C F
+red_ball.is_visible B F
+[…]
+Highway
+agent.{x,y}_position C F
+agent.x_velocity C T
+lane_existence.{upper,lower} B F
+vehicle_ahead.{x_position,x_velocity} C F/T
+vehicle_upper_lane.{x_position,x_velocity} C F/T
+vehicle_lower_lane.{x_position,x_velocity} C F/T
+Skiing
+[…]
+next_pole.y_position C F
+next_tree.{x,y}_position C F
+skier.x_position C F
+skier.x_velocity C T
+V-B2 VLM generating Feature Extraction code.
+For this step, we utilize claude-3.5-sonnet-20241022, one of the leading models in code generation.
 To address common challenges with VLM-generated code, such as non-functionality or runtime errors, we query the VLM 50 times using the same prompt (temperature = 0 for reproducibility) and select the most reliable feature extraction code through majority voting.
+Since code generation depends on segmentation quality $\mathcal{S}$, we further evaluate segmentation performance.
+Among the evaluated environments, Skiing presented the greatest difficulty for VLM-based segmentation labeling, and we therefore selected it for detailed analysis.
+On 100 sampled images, segmentation achieves high overall accuracy (Table III) but includes imperfect cases where the skier is partially or entirely missed (Fig.
+4).
+Despite such errors present in our in-context-learning examples (Fig.
+4), the generated code remains effective, demonstrating the robustness of our framework as long as a subset of segments is correctly labeled.
+(a) Skiing Exp. 1.
+(b) Skiing Exp. 2.
+(c) Skiing Exp. 3.
+Figure 4: Randomly selected segmentation examples we provide to the VLM for domain Skiing.
+Example 1 is good while 2 and 3 are imperfect.
+Example 2 identifies the skier as a partial skier and a noisy segment as skier.
+Example 3 identifies both the actual skier on the right and decoration elements at the bottom as partial skier.
 The generated codes are evaluated on a randomly collected dataset of 100 images and filtered based on the following criteria: (1) the code must compile and execute successfully and (2) it must produce diverse feature outputs rather than constant or repetitive values across the dataset.
 Candidates producing identical outputs are grouped into clusters, and the largest cluster is chosen as the final feature extraction code.
-Details are included in Appendix A.2.
-4.2.3 Extractor module performance analysis
-In Step 3 of our methodology, we address the inefficiencies of the VLM-generated feature extraction code for single-image frames, $\mathcal{G}_{s}$, by training an lightweight Extractor Module as its replacement.
-While $\mathcal{G}_{s}$ can be executed directly on the dataset, the processing pipeline is inherently non-batchable, as it handles image segmentations independently.
-This non-batchability severely impacts RL training, causing slowdowns of around 20000× compared to batchable operations (our CNN Extractor Module), as demonstrated in Table 3.
-From practical experience, this would translate to training times exceeding multiple weeks on an H40-level GPU for a single policy to converge.
-We constrain computing time to a reasonable range with a lightweight Extractor Module, resolving the slowdown caused by segmentation-based analysis in $\mathcal{G}_{s}$.
+TABLE III: Confusion ratios per feature concept
+Skier Pole Tree
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+V-B3 Extractor module performance analysis
+We train an CNN based Extractor Module to resolve the slowdown caused by segmentation-based analysis.
 For each environment, we train an CNN on 7,000 images collected from diverse trajectories, validate it on 2,000 images, and test it on 1,000 images.
-Compared to the VLM-generated ground truth, our models achieve high accuracy (typically >0.95) on most binary features, while MSE remains low (<0.005) for the majority of continuous features, with one exception where error is notably higher.
-Table 1 provides detailed performance statistics.
+As shown in Table IV, compared to the VLM-generated ground truth, our models achieve high accuracy (typically >0.95) on most binary features, while MSE remains low (<0.005) for the majority of continuous features, with one exception where error is notably higher.
 These results confirm that a lightweight vision model can serve as a reliable and computationally efficient replication for VLM generated extraction code $\mathcal{G}_{s}$.
-Table 1: Performance of the Extractor Module across environments, compared with ground truth from VLM-generated code.
+TABLE IV: Performance of the Extractor Module across environments, compared with ground truth from VLM-generated code.
 Continuous components are evaluated with MSE; boolean components are evaluated with accuracy.
 Dashes indicate vacant entries due to varying feature lengths and types across environments.
-Full feature descriptions are provided in Appendix A.2.2.
+Full feature descriptions are provided in Table II.
 Environment Feature Vector Indices (1–7)
@@
 Highway (Acc) 0.988 0.948
-4.3 Training results
+V-C Baselines
 We train ICT, our proposed interpretable and verifiable method that learns a tree structure with RL, alongside several baseline models across all environments.
 Below is an explanation of each baseline model used in the evaluation:
-  • CNN: An uninterpretable black-box model that maps raw visual inputs directly to the action space, trained with RL.
+  • CNN: An uninterpretable black-box model that maps raw visual inputs to the action space, trained with RL.
   • VLM-ICT: Our proposed interpretable and verifiable method, using an ICT to connect the semantic feature space with the action space, trained with RL.
   • VLM-ICT (Rand Init): Our proposed method, but initialized randomly and not trained with RL.
+  • VLM-EQL: An interpretable neural network that employs simple mathematical functions (e.g., square, cube, sine) to map semantic features to the action space [39].
   • VLM-MLP: An uninterpretable benchmark that connects a multi-layer perceptron between the semantic feature space and the action space.
@@
 (c) OpenAI Gym Highway
-Table 2: Runtime speedup on Skiing.
-Other domains in Appendix A.3.
-Batch VLM Code CNN Speedup
-8 19.056 0.001 13831.3x
-16 38.288 0.002 17325.6x
-32 76.212 0.004 19935.0x
-64 151.836 0.007 21692.9x
-128 308.468 0.014 22772.4x
-Figure 3: Results across three domains. Higher scores are better. Runtime comparison shows speedup of CNN-based extractor over VLM code.
+Figure 5: Results across three domains. Higher scores are better.
 In addition to standard learning-based baselines, we also explore whether a Vision-Language Model (VLM) can directly perform action selection via prompting, given access to a semantic feature space it previously constructed.
@@
   • VLM-DT (ChatGPT-4o-latest): A baseline where ChatGPT-4o generates a heuristic decision tree for action selection.
-  • VLM-DT (claude-3.5-sonnet): An equivalent baseline using Claude 3.5 Sonnet as the prompting engine.
+  • VLM-DT (claude-3.5-sonnet): An equivalent baseline querying Claude 3.5 Sonnet.
+V-D Train results
 We use Proximal Policy Optimization (PPO) as the reinforcement learning algorithm to train our policy [40], with a learning rate of 0.0005 and a discount factor of 0.99.
@@
 Results are shown in Fig.
-3.
-Across all three environments, VLM-ICT consistently outperforms other uninterpretable models (MLP) built on the semantic feature space, and demonstrates performance close to the CNN+RL black-box benchmark representing the upper limit of task performance.
-These consistent results highlight the robustness and effectiveness of our proposed method.
+5.
+Our results demonstrate the remarkable effectiveness of the proposed framework.
+As expected, the CNN policy achieves the highest performance across all three environments, benefiting from its status as a black-box model with orders of magnitude more parameters than the interpretable baselines.
+Notably, our proposed VLM-ICT method consistently reduces this gap.
+In Atari Skiing, VLM-ICT matches both the interpretable baseline (VLM-EQL) and the black-box policy (VLM-MLP) trained on the same features.
+In BabyAI and Highway, VLM-ICT outperforms the interpretable baseline (VLM-EQL) and reaches performance comparable to the black-box model (CNN, VLM-MLP).
+Furthermore, across all domains, our interpretable framework steadily outperforms decision trees generated directly by VLMs (VLM-DT ChatGPT and VLM-DT Claude), demonstrating that these are non-trivial control problems even when the semantic features provided are precisely task-relevant, with no redundant or irrelevant information.
 Building on a semantically meaningful feature set, our ICTs contains nodes directly corresponding to human-understandable concepts.
-While we do not evaluate interpretability directly, this approach is consistent with prior work where performance serves as a proxy for meaningful representations [14, 36].
+While we do not evaluate interpretability directly, this approach is consistent with prior work where performance serves as a proxy for meaningful representations [41].
 A previous human study [5] also supports the interpretability of such trees.
-We visualize the learned ICT in Appendix A.1, along with example cases demonstrating how the ICT yields plausible local explanations in example scenarios.
-5 Limitations and future works
-Our framework is yet limited by the assumption that SAM performs reasonably well across environments, which may hinder generalizability in domains with poor segmentation.
+V-E Takeaways
+Our results highlight followin achievements of the proposed iTRACE framework:
+  • Fast: By distilling VLM-generated feature extractors into a lightweight CNN, iTRACE avoids repeatedly querying the segmentation model at every timestep.
+This design yields several orders of magnitude speedup, making RL training practical at scale.
+  • Trainable: Both the feature extractor and the tree-based policy adapt effectively to new domains via gradient-based optimization.
+This enables pixels-to-action control and shows that interpretability does not come at the cost of flexibility.
+  • Interpretable: iTRACE provides transparency at every stage.
+The feature space is grounded in human-understandable concepts, and the learned Interpretable Control Tree ensures the resulting policy is verifiable.
+In summary iTRACE delivers a pixels-to-actions system that is simultaneously fast, trainable, and interpretable.
+VI Limitations and future works
+Our framework is limited by two factors: first, it assumes that SAM performs reasonably well across environments, which may hinder generalizability in domains with poor segmentation; second, it requires domain knowledge, which still involves human input to some extent and is left to be automated by foundation model queries or agentic exploration.
 Future work could explore RL-based backpropagation to fine-tune VLMs, allowing them to iteratively improve outputs based on agent feedback [42].
 This approach has the potential to enable VLMs to identify discrepancies between the underlying physics simulation and its visual representation, leading to improved feature extraction for downstream RL tasks.
-6 Conclusion
+VII Conclusion
 We introduce iTRACE, a novel framework for achieving automated semantic interpretability in RL.
-It automates feature space construction and extraction with zero-shot VLM querying, making it generalizable to unseen environments without human-in-the-loop.
-To overcome VLMs’ inability to generate valid decision policies, iTRACE leverages an ICT for RL training, obtaining transparent, human-understandable, and verifiable decision rules.
-The framework achieves performance on par with purely uninterpretable black-box models and black-box models built on the same feature space.
+iTRACE constructs interpretable feature spaces via zero-shot VLM querying and trains Interpretable Control Trees via RL to produce transparent, verifiable policies.
+It generalizes to unseen environments without human input and achieves performance comparable to black-box models.
 While all modern interpretable models, including tree-based approaches, face challenges in complex environments, iTRACE stands out among them as the first to achieve automated semantic interpretability in RL, marking a significant advancement that underscores its impact.
-Appendix A Appendix
-A.1 Appendix A: ICT Visualization
-Figure 4: Visualization of a highlighted decision path (outlined in red) in the BabyAI GoToRedBall environment, with the corresponding agent observation shown on the right.
-In this scenario, the agent’s state vector indicates it is close to the left wall (Left Wall = 1), not close to the front or right walls (Front Wall = 0, Right Wall = 0), and an object is detected (Obj Exist = 1) at ego-centric coordinates (Obj_x = 1, Obj_y = 4).
-At this sceneario, our learned ICT suggest the agent to move forward which aligns with humans intuition.
-Figure 5: Illustration of a highlighted decision path (outlined in red) in the BabyAI GoToRedBall environment.
-The corresponding agent observation is shown on the right.
-In this scenario, the agent’s state vector indicates it is not close to any wall (Left Wall = 0, Front Wall = 0, Right Wall = 0), and no object is detected (Obj Exist = 0).
-The default object location is set to Obj_x = 0, Obj_y = –5.
-Since the object is not presented in agent’s view, our learned ICT suggests the agent to turn left to search for the object.
-A.2 Appendix B: VLM Prompting Ablation
-A.2.1 Segmentation Labeling
-Among all evaluated environments, Skiing posed the greatest challenge for LLM-based segmentation labeling.
-To reduce expensive API usage while showcasing a representative case, we center our analysis on this environment.
-As shown in Figure 6, segmentation quality is imperfect: the skier is sometimes only partially labeled or missed entirely due to large appearance variations across frames.
-BabyAI and Highway environments contain many repetitive segments that occupy a large portion of each image.
-To stay within the VLM’s context window token limit, we filter out strictly identical segments using simple heuristics such as aspect ratio and area.
-This assumes that highly frequent, repetitive elements are less critical for feature extraction.
-This limitation can be mitigated as LLM context capacity increases.
-Once longer context windows are supported, all segments can be included, and irrelevant decorative elements can still be correctly labeled and filtered, as shown qualitatively in our examples.
-(a) Actual skier is labeled correctly, but decoration elements at the bottom are labeled as partial skier.
-(b) Noisy segment is identified as skier, while the actual skier is labeled partial.
-(c) Skier identified as decoration.
-Figure 6: Examples of imperfect segmentations in the Skiing environment.
-Despite errors in example segmentation, our method successfully retrieves effective extraction functions for key features across environments, and the resulting code generalizes well across frames.
-This demonstrates that our pipeline is robust to segmentation errors: as long as a subset of segments is labeled correctly, the system can proceed to generate reliable feature extraction logic.
-Reproducibility.
-To support reproducibility, we provide the specific examples and their LLM segmentations for all three environments. Some of them are imperfect.
-(a) Skiing Example 1.
-(b) Skiing Example 2.
-(c) Skiing Example 3.
-Figure 7: Randomly selected segmentation examples we provide to the VLM.
-Example 1 is good while 2 and 3 are imperfect.
-Example 2 identifies the skier as a partial skier and a noisy segment as skier.
-Example 3 identifies both the actual skier on the right and decoration elements at the bottom as partial skier.
-(a) BabyAI example 1.
-(b) BabyAI example 2.
-(c) BabyAI example 3.
-Figure 8: Randomly selected segmentation examples we provide to the VLM. Some segmentation is imperfect. For example, the green box is not correctly identified in example 2.
-(a) Highway example 1.
-(b) Highway example 2.
-(c) Highway example 3.
-Figure 9: Randomly selected segmentation examples we provide to the VLM for Highway. Segmentation for this environment is mostly perfect.
-A.2.2 Feature Dictionary Extraction
-Here we provide the complete list of features that the VLM decided for each environment in Table. 3, 4, 5 .
-Table 3: Feature schema for BabyAI RedBall environment.
-Object Feature Type Requires Temporal Context
-red_ball x_position Continuous No
-y_position Continuous No
-is_visible Categorical No
-obstacle_forward is_blocked Categorical No
-obstacle_left is_blocked Categorical No
-obstacle_right is_blocked Categorical No
-Table 4: Feature schema for Highway environment.
-Object Feature Type Requires Temporal Context
-agent x_position Continuous No
-y_position Continuous No
-x_velocity Continuous Yes
-lane_existence upper_lane_exists Ordinal No
-lower_lane_exists Ordinal No
-vehicle_ahead_same_lane x_position Continuous No
-x_velocity Continuous Yes
-vehicle_adjacent_upper_lane x_position Continuous No
-x_velocity Continuous Yes
-vehicle_adjacent_lower_lane x_position Continuous No
-x_velocity Continuous Yes
-Table 5: Feature schema for Skiing environment.
-Object Feature Type Requires Temporal Context
-next_gate x_position_left_pole Continuous No
-x_position_right_pole Continuous No
-y_position Continuous No
-closest_tree x_position Continuous No
-y_position Continuous No
-skier x_position Continuous No
-x_velocity Continuous Yes
-A.2.3 Extraction Function Retrieval
-Table 6: Prompting statistics for VLM-generated function 1 across 50 trials.
-“Code failure” indicates code that cannot import or run.
-“Excluded as constant" refers to code that generates invariant outputs across diverse samples and is therefore excluded.
-“Majority selection coverage” represents the proportion of code yielding identical outputs in the majority voting process.
-Step 1 refers to the process of generating $\mathcal{G}_{s}$, while Step 2 refers to the process of generating $\mathcal{G}_{m}$.
-Env. Code Failure Rate (%) Excluded as Constant (%) Majority Selection Coverage (%)
-Skiing 8 0 72
-BabyAI 8 74 4
-Highway 0 90 4
-To address the unreliability of VLM-generated code, particularly hallucinations and execution failures, we implement a selection procedure to identify robust feature extraction functions.
-For each of the three environments, we prompt the VLM to generate 50 candidate functions, evaluate them on a dataset of 1000 images, and filter out those that either fail to run or produce invariant outputs across all samples.
-From the remaining valid functions, we apply majority voting based on output similarity and select the function whose outputs most frequently align with others.
-This strategy assumes that the dataset is sufficiently diverse such that no single feature remains constant across all frames.
-Table 6 presents statistics from this process, including code failure rates, exclusions due to constant outputs, and majority selection coverage.
-While the majority selection coverage appears relatively low, this is primarily due to inconsistent normalization of features representing relative distances.
-In contrast, features encoding object positions tend to be represented more consistently across functions.
-A.3 Appendix C: CNN Performance Analysis
-A.3.1 CNN structure for the Skiing environment.
-The SkiingCNN model is a general-purpose convolutional network designed for multiple environments.
-For the skiing environment specifically, it outputs a 6-dimensional feature vector per frame, capturing semantic information such as skier position, gate pole positions, and nearby trees.
-  • Input: A $3\times 56\times 76$ RGB segmentation image, with distinct segments corresponding to skier, gates, and trees.
-  • Environment Routing: At initialization, the network identifies its task from an environment string (e.g., env_n = "skiing").
-This string determines the output dimensionality; for skiing, it is fixed to 6.
-The model uses a compact, task-specific CNN with the following layers:
-  • Conv2d(3 $\rightarrow$ 32, kernel=5) $\rightarrow$ ReLU $\rightarrow$ MaxPool2d(2) $\rightarrow$ [32, 25, 19]
-  • Conv2d(32 $\rightarrow$ 64, kernel=3) $\rightarrow$ ReLU $\rightarrow$ MaxPool2d(2) $\rightarrow$ [64, 5, 4]
-  • Conv2d(64 $\rightarrow$ 64, kernel=3) $\rightarrow$ ReLU $\rightarrow$ [64, 3, 2]
-This yields a low-resolution spatial encoding with 64 channels and a flattened size of 384, computed dynamically during initialization.
-MLP Output Head. The fully connected layers process the flattened feature vector to produce semantic outputs:
-  • Linear(flat_size $\rightarrow$ 512) $\rightarrow$ ReLU $\rightarrow$ Dropout(0.4)
-  • Linear(512 $\rightarrow$ 6) for skiing outputs
-Output Format.
-The final 6-dimensional output contains semantically meaningful quantities related to skiing dynamics.
-Unlike in some other environments (e.g., RedBall), no post-hoc sigmoid activation is applied, as all skiing outputs are treated as continuous and directly regressed.
-A.3.2 CNN structure for the BabyAI-GoToRedBall environment.
-For this part, we use a combination of two CNN networks, BinaryDetectCNN and BallDetectionCNN.
-The BinaryDetectCNN is a multi-task convolutional model designed to predict (1) binary obstacle existence in the forward, left, and right directions, and (2) the existence of a red ball in the frame.
-The model uses shared convolutional features and two dedicated branches:
-  • Shared Feature Extractor: Three convolutional layers with channels [3 $\rightarrow$ 32 $\rightarrow$ 64 $\rightarrow$ 128], each followed by batch normalization and ReLU.
-The feature map is downsampled with stride-2 convolutions instead of pooling for better spatial preservation.
-  • Block Detection Head: A dedicated encoder (1 conv layer) followed by adaptive average pooling and a two-layer MLP [128 $\rightarrow$ 256 $\rightarrow$ 3] for predicting obstacle presence in three directions.
-The output is passed through sigmoid activation.
-  • Ball Existence Head: A parallel encoder with the same structure, yielding a single sigmoid output indicating the presence of a red ball.
-  • Output Format: The final output is a 6-dimensional vector: 3 binary obstacle indicators, 2 position placeholders, and 1 binary ball existence score.
-The placeholder values are used to maintain a consistent output shape and are overwritten when positional regression is available.
-The BallDetectionCNN is a separate network responsible for regressing the $(x,y)$ position of the red ball, but only if the ball is detected by BinaryDetectCNN.
-  • Convolutional Backbone: A four-block CNN with increasing channels [3 $\rightarrow$ 32 $\rightarrow$ 64 $\rightarrow$ 128 $\rightarrow$ 256], each block consisting of convolution, batch normalization, ReLU, and $2\times 2$ max pooling.
-  • Adaptive Pooling: The output is pooled to a $1\times 1$ spatial map to support arbitrary input sizes.
-  • Regression Head: A three-layer MLP [256 $\rightarrow$ 128 $\rightarrow$ 64 $\rightarrow$ 2] outputs the normalized coordinates of the ball.
-  • Inference Logic: The regressor is only invoked if the ball existence score from BinaryDetectCNN exceeds 0.5.
-The image is resized to $64\times 64$ and normalized before being passed to the regressor.
-At inference time, both models are used jointly to generate a compact 6-dimensional feature vector:
-[left_blocked, front_blocked, right_blocked, ball_exists, ball_x, ball_y]
-If the ball is not detected, the position fields are filled with default placeholder values (e.g., $x=0$, $y=-5$).
-This conditional two-stage pipeline enables efficient and accurate feature extraction without overfitting to spatial priors or dense supervision.
-A.3.3 CNN structure for the Highway environment.
-To improve regression accuracy for specific vehicle positions, we deploy three additional networks, each a specialized regressor based on ResNet-18. These models handle:
-  1. 1.
-    Vehicle ahead in the same lane
-  2. 2.
-    Vehicle in the adjacent upper lane
-  3. 3.
-    Vehicle in the adjacent lower lane
-Each HighwayCNN model follows the structure below:
-  • Backbone: A ResNet-18 pretrained on ImageNet. All layers are initially frozen, with the last three residual blocks selectively unfrozen for fine-tuning.
-  • Custom Head: The original fully connected layer is replaced with a regression head: [512 $\rightarrow$ 256 $\rightarrow$ 128 $\rightarrow$ 1], with batch normalization, ReLU activations, and dropout (0.3, 0.2).
-  • Output: A single scalar value representing the $x$-position of the target vehicle.
-At inference time, the models are used jointly to construct a 7-dimensional feature vector per frame:
-[lower_lane, upper_lane, agent_x, agent_y, vehicle_ahead_x, vehicle_upper_x, vehicle_lower_x]
-The main CNN provides lane existence and agent position.
-The three auxiliary regressors are queried independently to estimate the x-coordinates of nearby vehicles in the ego, upper, and lower lanes.
-This multi-stage pipeline allows for flexible specialization across sub-tasks while maintaining a compact, semantically meaningful feature representation.
-A.3.4 Additional Runtime Analysis
-We used a more complicated CNN structure which slows down CNN exponentially in those two domains, but the CNN computational time is still significantly non-trivially shorter than the actual computing time.
-Table 7: Runtime speedup on BabyAI
-Batch VLM Code CNN Speedup
-8 42.435 (s) 0.037 (s) 1132.93x
-16 97.351 (s) 0.042 (s) 2303.66x
-32 187.182 (s) 0.097 (s) 1924.26x
-64 376.923 (s) 0.112 (s) 3361.79x
-128 750.582 (s) 0.342 (s) 2196.65x
-Table 8: Runtime speedup on Highway.
-Batch VLM Code CNN Speedup
-8 42.421 (s) 0.025 (s) 1683.08x
-16 88.326 (s) 0.040 (s) 2207.01x
-32 172.973 (s) 0.084 (s) 2047.23x
-64 347.060 (s) 0.186 (s) 1868.32x
-128 696.099 (s) 0.328 (s) 2122.57x
-A.4 Appendix D: Prompts
-We employ Chain-of-Thought prompting combined with structurally formatted Markdown outputs to guide the model toward producing target results.
-The complete prompts used for segmentation labeling, feature extraction, Function 1 generation, and Function 2 generation are provided below.
-A.4.1 Segmentation Prompt
-You are tasked with interpreting every segment from a segmented image and labeling it with what it appears to represent.
-You will be given a single image called image.png, along with a segmentation file masks.json generated by Segment Anything.
-This file contains a list of all visual regions detected in the image.
-Each segment includes attributes such as bounding box, area, dominant color, and other visual metadata.
-Your goal is to assign a predicted_object label to every segment in this list, whether the segment depicts a featured object, a decorative element, or a fragment of a larger object.
-Note that the color attribute is given by the most dominant color in the segment. If the segment is colorful, the dominant is not reliable and you should not depend on it.
-You will first receive a feature dictionary specifying which features are to be derived from the image.
-Begin by examining this dictionary to determine which object types are necessary in order to compute the values of those features.
-You may ignore any features marked with "requires_temporal_context": True, as they are not relevant for single-frame analysis.
-Based on the remaining features, deduce what kinds of objects must be identified to extract those values.
-These are your featured objects, and you must list them explicitly before proceeding to segment analysis.
-The feature dictionary is defined here: <insert-features>
-Domain Knowledge: <insert-domain-knowledge>
-Do not assume that visually large, centered, or high-contrast segments are featured objects unless it aligns with what you see from the image I provided.
-A segment should only be labeled as a featured object if its shape, size, position, and color closely match a plausible instance of that object in the image.
-The presence or prominence of an object in the image is not evidence of its relevance unless it is explicitly needed for feature extraction.
-If the feature dictionary requires you to distinguish objects by their relative positions (e.g., “ahead”, “behind”, “left of”, “right of”, or “same group”), you must assign them a general segmentation label, and leave any spatial analysis alone.
-A separate stage will handle positional analysis later, so you don’t need to worry about it here.
-Once you have identified the featured objects, begin a two-pass labeling process over the segmentation data.
-For the first pass, analyze each segment independently.
-Use only the attributes provided in the masks.json file—such as bounding box, area, aspect ratio, dominant color, and relative position—and your visual observation of the image.
-Base your predictions strictly on observed features and metadata.
-Do not infer identities from prior knowledge or gameplay expectations.
-Assign each segment a predicted_object label that reflects what it is most likely to represent.
-If a segment matches one of the featured objects you previously identified, label it with that object’s name.
-If a segment appears to be a visual decoration (such as background, UI overlay, or border), label it explicitly as such.
-If a segment clearly represents only a part of a larger object, assign it a label indicating it is a fragment (e.g., "fragment" or "partial_<object>").
-If you cannot confidently classify a segment, label it as "ambiguous".
-After this first pass, review your labeling decisions in context.
-Consider whether your predictions make sense collectively.
-For example, ask whether any object that should only appear once is represented by multiple segments, whether any key object is missing entirely, or whether the traits of any labeled segments contradict the visual evidence from the image.
-If you identify inconsistencies, refine your labels accordingly to make the overall scene interpretation more coherent.
-The second pass should resolve overlaps, mismatches, or missing predictions that become evident only in context.
-Your output must return the list of segments from the original masks.json, preserving all original fields while adding a new "predicted_object" field to each segment dictionary.
-Every segment must be labeled.
-No segment may be left without a prediction, even if it is decorative, irrelevant, or ambiguous.
-Segmentation data derived from the image is here: <insert-segment-knowledge>
-Begin your reasoning inside the Reasoning section, where you strictly follow the thinking process:
-  1. 1.
-    Ignore the segmentation and look at the image. Describe all elements you observe, including all decorational ones.
-  2. 2.
-    Look at the feature dictionary. Deduce the featured objects required by the feature dictionary.
-  3. 3.
-    Analyze each segment independently and assign them with a prediction label following the instructions for the first pass.
-  4. 4.
-    Conduct the second pass.
-Finally, return the updated segment list with predicted_object annotations inside the ## Final Answer section.
-Your final answer should follow the same format with the input, strictly copying the values while adding your predictions:
-    "area": area,
-    "bbox": [x, y, w, h],
-    "predicted_iou": predicted_iou,
-    "point_coords": [
-      [x, y],
-    "stability_score": stability_score,
-    "crop_box": [x, y, w, h],
-    "dominant_color": [r, g, b],
-    "predicted_object": object_name
-Your output should follow this structure:
-## Reasoning
-…all steps…
-## Final Answer
-  { "area": ..., "predicted_object": "your prediction", ... },
-A.4.2 Feature Extraction Prompt
-Instructions: You are given a series of image frames from a specific environment, and your task is to determine the features essential for making the immediate decision.
-Focus on extracting the positional data of key objects and any associated features that can aid in choosing the immediate movement.
-Ignore any elements that are visible but do not affect the agent’s ability to choose its immediate movement.
-This includes elements whose behavior is entirely deterministic once triggered or which cannot be altered by future actions.
-You should also ignore anything that cannot be directly inferred from the image or requires assumptions beyond visual evidence.
-If certain properties of an object (such as its type, color, or exact shape) do not influence what the agent will do next, you are encouraged to abstract them away.
-The goal is to capture only the information needed to distinguish between available actions.
-Extraneous visual or semantic detail should be excluded, even if it is visible in the frame.
-You may also consider changes in object positions across time (e.g., motion trends) when necessary.
-In some environments, both horizontal and vertical positions/velocities of key elements across frames are informative.
-They should be always separated by direction and axis (e.g., x_position, y_position, x_velocity, y_velocity).
-Positional or motion features must be separated into explicit horizontal (x) and vertical (y) components whenever relevant.
-If either direction (horizontal/vertical) doesn’t matter, exclude it; otherwise, always include both horizontal and vertical direction.
-If movement trends are not relevant to decision-making in a given case, they should be omitted entirely.
-Not all information visible in the image is essential for decision-making. Disregard anything that isn’t.
-Important: Some environments may contain multiple instances of the same type of object.
-You should not assume a fixed or singular representation for such objects unless explicitly outlined in the environment description.
-When considering non-goal objects that could impact the movement (obstacles, enemies, etc.), avoid modeling the entire map or distant entities.
-Instead, identify whether anything in the agent’s immediately reachable space would prevent it from either choosing or executing any of its primitive actions.
-Most commonly, this will be the closest instance (obstacles, enemies, etc.), as that is most likely to impact immediate behavior.
-You must not include other instances unless their effect is simultaneous and unavoidable.
-You should also explain your selection criterion under Final Description.
-When identifying environmental elements that constrain the agent’s immediate behavior, consider whether anything in the agent’s directly actionable space—that is, any region reachable by a single primitive action—would limit or prevent the choice of that action.
-For instance, an obstacle to the side may not block movement directly but could influence whether turning should be chosen as the next immediate action.
-Do not include distant or visually salient elements unless their effect is immediate and unavoidable.
-You should also abstract away any properties (such as type, color, or identity) that do not influence the outcome of the agent’s next action.
-You should not ground your analysis in the specific frames you’re shown, except if things stay consistent across them.
-These are only examples to help illustrate the structure of possible observations.
-Your job is not to analyze the example frames individually, but to construct a feature dictionary that would apply across frames and across environments sharing the same interaction mechanism.
-This means the actual position or visibility of the target object in the examples should not influence your output.
-Domain knowledge: <insert-domain-knowledge>
-Your answer should be structured as:
-  "element 1": [
-        "attribute 1": {
-            "values": ["option1", "option2", ...],
-            "type": "ordinal",
-            "requires_temporal_context": False
-        "attribute 2": {
-            "values": x,
-            "type": "continuous",
-            "requires_temporal_context": False
-        "attribute 1": {
-            "values": x,
-            "type": "continuous",
-            "requires_temporal_context": True
-        "attribute 2": {
-            "values": x,
-            "type": "continuous",
-            "requires_temporal_context": False
-  "element 2": {
-        "attribute 1": {
-            "values": ["option1", "option2", ...],
-            "type": "ordinal",
-            "requires_temporal_context": True
-Each element should represent an environmental object or factor that could influence the agent’s immediate decision.
-If multiple elements of the same type are relevant, they should be provided as a list.
-Each attribute describes a property of that element that could affect the decision-making process.
-Please do not make up or fill in specific values like 60, 100, or -5 in the "values" field.
-That field should remain blank or placeholder-like (e.g., "values": "continuous") to indicate the type and temporal context of each attribute.
-Your output should follow this structure:
-## Reasoning
-…all steps…
-## Final Answer
-(stricly follow the format specified above)
-## Description
-Describe the visual traits of the featured objects and the selection mechanism from a list of identical objects (if applicable).
-This should help a third-party reader understand what your features refer to, not explain your reasoning process.
-A.4.3 Function 1 Generation Prompt
-You are tasked to write a python function to retrieve values of the specified features from an image in a consistent environment.
-This function should be able to generalize across frames.
-Describe all positions by ratio in the image x y axis.
-Your code should take in one single image frame (image.png) along with a json dictionary (masks.json) given by Segment Anything on the image.
-Name the function you write ‘extract_features(img_dir, json_dir)‘.
-Note that the segments in the JSON dictionary contain every possible element in the image frame, many of which are irrelevant or decorative.
-Thus, you must include logic to exclude such elements and only retain the segments that are visually meaningful for the defined features.
-Each entry in masks.json corresponds to a visually distinct segment in the image.
-Note that these segments include both relevant objects (e.g.
-boxers) and irrelevant or decorative elements (e.g.
-background, UI, borders, large static surfaces, or visually dominant non-interactive areas).
-Note that this segmentation file provides a list instead of a dictionary.
-I will provide you with some example images and their segmentation data printed in text form, but note that the function you write will deal with a completely different image and its different segmentation data.
-The example images are embedded with visualizations of the segmentations to give an idea of what the targets and the noises will look like, but the actual image your function deals with won’t be.
-In the example segmentation data, each segment includes a "predicted_object" field which labels what the segment likely represents.
-This field exists only for the sake of demonstration—to help you learn what kinds of attributes distinguish key objects from background clutter.
-The "predicted_object" field will NOT be present in the actual masks.json file that your function processes.
-Therefore, your final code must NOT rely on this field in any way.
-Instead, use it in the examples purely as reference to derive the selection logic.
-However, you must study the "predicted_object" labels in the examples carefully to identify consistent, distinguishing traits for each object type you intend to detect.
-Your selection logic should be explicitly based on those shared traits—such as area, color composition, aspect ratio, and position—that reliably differentiate the intended objects from irrelevant segments like UI, background, or decorative fragments.
-Do not use overly broad filters unless you have confirmed—through all provided examples—that such filters exclude all non-target segments without failure.
-Important: Raw segmentation masks are not provided. You must instead analyze the image directly using the information from the image pixels as a visual clue if needed.
-Some features may require information that goes beyond what’s available in the segments alone.
-While segments provide a reliable reference for object-level statistics, they may not capture everything—such as irregular shapes that aren’t well represented by bounding boxes or aspect ratios, or subtle lighting conditions where average color is insufficient.
-In such cases, you should refer directly to the image rather than relying solely on the segmentation.
-Segmentation data derived from the example images: <insert-segment-knowledge>
-You should determine the feature objects you want to search for based on the domain knowledge and target feature dictionary I give you, and think about how to select them from the segmentation data.
-Domain knowledge: <insert-domain-knowledge>
-Features that the function is tasked to derive: <insert-features>
-where each individual dictionary item represents an environmental object or factor that could influence the agent’s next decision.
-If you identified multiple objects of the same type, they should be provided as a list in the dictionary item as specified by examples in the given feature dictionary.
-For the attributes of each object, there will be a "requires_temporal_context" field which is either True or False (boolean).
-If it’s False, neglect the attribute and don’t return anything.
-Only if it’s True, analyze the image and the Segment Anything masks to decide a value.
-For the attributes of each object, there will be a "type" field which would be either "continuous" or "ordinal".
-If marked "continuous", your code should derive a number value.
-If marked "ordinal", your code should return one of the specified ordinal values as a string.
-Additional clarifications about the features can be found here.
-You should use it either (1) as supplementary information about the object’s visual traits to help your selection, or (2) as strict criteria that must be enforced when selecting a type of objects, if any specific requirements are stated.
-<insert-explanation>
-Your code should return a dictionary keeping all the original dictionary items, while only including attribute values and excluding the "type" and "requires_temporal_context" fields.
-Your code should be able to run as it is, without requiring further information or edit.
-Note that every featured object exists in the example image, but not necessarily in all images the function deals with, so return None if you can’t find the featured object.
-Segment Anything already provides well-grouped regions for each object — do not merge or cluster multiple segments.
-If the featured object is supposed to appear only once, you must select exactly one segment to represent it.
-Note that the color attribute is given by the most dominant color in the segment. If the segment is colorful, the dominant is not reliable and you should not depend on it.
-Objects will vary across frames.
-You are exposed to only a limited sample of image frames, so it’s advised to keep the selection criteria loose as long as you can verify it doesn’t lead to misselecting any decorative elements, UI backgrounds or fragments.
-If the selection criteria involves spatial relationships and occupancy between general objects, you must NOT use the object’s bounding box or center to make this decision.
-Instead, you should use pixel-wise content in the image region corresponding to the spatial location and determine the featured value based on its actual pixel content.
-Regions that are not occupied tend to exhibit consistent visual characteristics—such as uniform texture or stable color patterns—while occupied regions vary more widely in appearance.
-You should resort to the domain knowledge for additional information and visual cues of representative patterns observed in known examples.
-If the selection criteria involves spatial relationships between fully-filled objects, you should account for the object’s full bounding box, not just its center.
-However, this is only allowed when the goal is to estimate a continuous relative offset, not binary occupancy.
-An object may influence or obstruct another’s path if any part of its area extends into the relevant region, even if its center does not.
-If the feature dictionary involves discrete categories of relative positions, you must NOT divide the image into uniform bands or grids along the height, as the viewpoint may shift across frames.
-Instead, you should define fixed-width zones based on relative spatial relationships of visual landmarks or reference objects visible in the current frame.
-Additional Instruction: If a featured object is represented by multiple segments of the same type, it is acceptable—and sometimes necessary—to first retrieve a related value that is not included in the final feature dictionary to aid in the selection process.
-Strictly follow the defined selection criteria, and do not hesitate to compute temporary or auxiliary values if they are essential for choosing the correct segment.
-However, these intermediate values should not be included in the final dictionary output unless they are explicitly required.
-Your thinking should follow these steps.
-– Look at the segmentations for each example image and the "predicted_object" field (only available in the examples), assuming that the prediction is correct.
-Pay attention to the attributes of featured objects – what traits can be used to select them from the segmentation json?
-– Discuss and outline your approach for separating decorative elements from the specified key objects.
-Would any of the decorative elements, especially borders, UI, background, fall into your selection category?
-How did you ensure that the segment you selected is the actual item?
-If the item is supposed to be singular instead of a list, how did you ensure only one segmentation would qualify under your criteria as the actual item?
-– Describe your object selection logic and your plan of writing a python function that detects the positions of the objects of interest detailedly.
-If some segmentations are clearly decorative or part of the background (e.g., one of their statistics is noticeably different), it’s recommended to filter them out before selecting the actual object to avoid errors.
-Again, since objects will vary across frames, you should keep the object selection criteria as loose as possible.
-– Simulate your planned selection logic on the full segmentation data, one by one, including not just featured objects but also irrelevant segments, UI backgrounds, and partial fragments.
-You must verify that your method consistently identifies the correct segments and does not mistakenly include irrelevant ones.
-Focus on the numerical values on the first segmentation example I gave you.
-If your strategy fails (e.g., selects a background or UI element), refine it before proceeding.
-You must prove to yourself that the approach works on the provided examples.
-You must validate that your approach works without relying on "predicted_object" in actual usage.
-## Reasoning
-## Code
-A.4.4 Function 2 Generation Prompt
-You are tasked to write a python function to retrieve values of the specified features from two chronological images in a consistent environment, where object traits remain stable.
-The timestamp between the images will always stay consistent.
-Your code should take in two image frames (png) along with their associated json dictionaries (json) given by Segment Anything on the images.
-Your function should be named after ‘def extract_features_two_frames(img1_dir, json1_dir, img2_dir, json2_dir)‘.
-Domain knowledge: <insert-domain-knowledge>
-Features that the function is tasked to derive: <insert-features>
-where each individual dictionary item represents an environmental object or factor that could influence the agent’s next decision.
-If you identified multiple objects of the same type, they should be provided as a list in the dictionary item as specified by examples in the given feature dictionary.
-For the attributes of each object, there will be a "type" field which would be either "continuous" or "ordinal".
-If marked "continuous", your code should derive a number value.
-If marked "ordinal", your code should return one of the specified ordinal values as a string.
-For the attributes of each object, there will be a "requires_temporal_context" field which is either True or False (boolean).
-This field indicates whether the attribute value requires two images frames or one single image frame to determine.
-In a previous stage, you already wrote a python function that takes in one single image frame and its associated segmentation information (input: image.png, masks.json) and returns a dictionary keeping all the original dictionary items, only including attribute values and excluding the "type" and "requires_temporal_context" fields.
-For each attribute, the function determines an actual value if "requires_temporal_context" is False, and inserts None if it’s True.
-The function is:
-<insert-function1>
-In this step, you should write a python code that utilizes the previous function for single image.
-Your code should take in two image frames (image1.png, image2.png) along with their associated json dictionaries (masks1.json, masks2.json), apply the single image function to them separately, and then return a dictionary keeping all the original dictionary items, only including attribute values and excluding the "type" and "requires_temporal_context" fields.
-Your code should be able to run as it is, without requiring further information or edit.
-Note that every featured object exists in the example image, but not necessarily in all images the function deals with, so account for that in your code logic.
-For each attribute, your code should return the value of running the previous function on the second image and completely ignore the value from the first image if "requires_temporal_context" is False.
-If "requires_temporal_context" is True, you should either calculate from the results of running the previous function on both images, or if that’s still not sufficient for the attribute, analyze the image and segmentation information from both images to retrieve a value.
-If a velocity or any other attribute requires using two values, and either value is None, then your code must return None for that attribute.
-Describe all positions by ratio in the image x y axis.
-I will provide you with three example images and their segmentation data printed in text form, but note that the function you write will deal with different images and different segmentation data.
-Segmentation data derived from the example images: <insert-segment-knowledge>
-Begin by explaining your reasoning step-by-step.
-After that, write the required python function.
-Double check that if an attribute requires using two values, either value being None would not affect your output.
-## Reasoning
-## Code
-A.4.5 Domain knowledge.
-The domain knowledge of each environment can be found below respectively.
-Skiing
-You control a skier who moves strictly downhill and can only shift sideways (left or right) by tilting your skis in that direction.
-The game uses a relative camera view: the skier remains approximately fixed near the top-center of the image (though not at the exact top), while the camera scrolls with the skier’s downhill motion.
-However, the camera does not follow the skier’s lateral movement, so their sideways position within the frame varies.
-In image space, the downhill direction corresponds to moving toward the bottom of the screen.
-The skier cannot move upward, so you must only consider gates and trees that appear below the skier in the image.
-The skier is primarily moving downhills towards the bottom of the image, which is also its primary goal.
-You must ignore gates and trees above the skier since they are no longer reachable.
-The goal is to pass through all gates as quickly as possible while avoiding trees below the skier.
-A gate consists of the horizontal space between two poles positioned at the same latitude.
-The skier must pass between both poles of the same gate to successfully clear it.
-You are penalized five seconds for each gate you miss.
-If you hit a gate or a tree, your skier will jump back up briefly before continuing downhill.
-You should avoid extreme sideways movement which is indicated by a noticeable drop in skier speed.
-In such situations, you must actively tilt the skis to regain speed and keep moving.
-Otherwise, the skier will gradually get stuck.
-Actions Skiing has the action space of Discrete(3) with the table below listing the meaning of each action’s meanings.
-Value | Meaning
-    0 | NOOP
-    1 | RIGHT
-    2 | LEFT
-Highway
-You control a car that drives forward along a multi-lane highway and can change lanes or adjust its speed to navigate traffic.
-The lanes appear as vertical bands (spanning top to bottom), and vehicles move horizontally across them from left to right.
-The road contains multiple parallel lanes arranged vertically, with some lanes located closer to the upper edge and others closer to the lower edge.
-Vehicles travel along these lanes and may appear in the same, adjacent upper, or adjacent lower lanes relative to the agent.
-Not every lane is directly next to an edge, and the number of visible lanes will vary depending on your position.
-Your task is to drive efficiently by anticipating traffic behavior and adjusting your actions to preserve forward momentum.
-You can switch lanes upward or downward, as long as such a lane exists and is not blocked.
-Other cars on the road may accelerate, decelerate, or change lanes, making them dynamic obstacles.
-Adjacent or closely ahead vehicles in the immediate upper or lower lanes present the most immediate risk during a lane change and should be closely monitored for lateral overlap and motion trends.
-The road is defined by visible lane markers: dashed lines separate adjacent lanes, while solid lines mark the outer boundaries.
-If the vehicle is next to a solid line, it means there is no adjacent lane available for switching in that direction.
-Veering beyond the highway boundaries would result in an immediate crash.
-The distance between lanes is uniform, meaning the gap between any two neighboring lane markers represents the standard lane interval.
-The width (height) of a car is notably smaller than the width (height) of an entire lane, and thus should not be used as a measure of lane width.
-The vertical distance between lane markers reliably measures lane width.
-Lane marker distance can be confidently used to determine if a car is in an adjacent lane or in further lanes.
-If you want lane width, you must calculate it as the distance between lane markers. You must NOT assert it from the segmentations.
-Collisions result in episode termination and must be avoided.
-You are rewarded for reaching and maintaining higher speeds without incident.
-Smooth driving and intelligent maneuvering through dense traffic are key to success in this environment.
-Actions Driving has the action space of Discrete(5) with the table below listing each action’s meaning.
-Value | Meaning
-0 | IDLE
-1 | FASTER
-2 | SLOWER
-3 | SWITCH_UPPER
-4 | SWITCH_LOWER
-BabyAI
-In this BabyAI GoToRedBall environment, the world is an enclosed grid.
-The agent is dropped into a new configuration each episode, surrounded by a few simple objects of different colors.
-Your goal is to go to the red ball.
-A complete list of all possible items that will occur in the environment: color type
-color is the color of the box. Can be “red”, “green”, “blue”, “purple”, “yellow” or “grey”.
-type is the type of the object. Can be “ball”, “box” or “key”.
-There will be one and only one red ball. The red ball may be present anywhere in the environment, and it may or may not be visible in a given frame.
-The agent appears as a red triangle positioned at the center of the bottom row, always facing upward.
-The edge of the visible world is flanked by a uniform grey strip, which separates a lighter interior grid from a darker exterior region.
-No objects appear beyond this strip.
-There are no doors in this environment.
-The observation is ego-centric: the agent is always placed at the center of the bottom row in the image.
-It can see all tiles directly in front of it, as well as up to three tiles to each side—limited only by the surrounding walls.
-The image consistently covers a 7×7 grid area, centered horizontally around the agent’s position.
-You need to control the agent to navigate to the red ball.
-The agent cannot move onto a grid cell that contains an object or a wall.
-As a result, any attempt to enter an occupied cell or pass through a wall is invalid and should be avoided.
-Any attempt to turn towards an occupied cell should also be avoided unless the agent has long-term intentions (e.g.
-making an U-turn to go back).
-BabyAI uses an action space of Discrete(7), but only the first three actions are used:
-Value | Meaning
-   0   | Turn left
-   1   | Turn right
-   2   | Move forward
-   3-6 | Unused
-A.5 Appendix E: VLM Hallucination Analysis
-Intuitively, one might also consider using VLMs to label features for a dataset of images.
-While VLMs may not provide precise positional values, coarse ordinal spatial descriptors, such as "left," "far left," or "right", can be sufficient for environments like skiing.
-However, our experiments demonstrate that this approach is fundamentally limited.
-Most VLMs visual backbones, such as CLIP or SigLIP, encode images by dividing them into fixed-size patches (e.g., 32×32 pixels).
-Each patch is linearly embedded and processed through transformer layers to model contextual relationships across patches.
-However, this patch-based architecture struggles to maintain accurate spatial relationships between objects when multiple objects fall within the same patch.
-This is a critical failure for feature extraction, as relevant features can appear anywhere in the image, and distinguishing between multiple objects within a single patch becomes impossible, even when grid markers are added to disambiguate spatial zones.
-To illustrate this issue, we conducted a simple experiment (Figure 10). When prompted with “What do you see? Describe tile by tile,” chatgpt-4o-latest responded:
-  • First tree (smaller): around x2y15 (bottom-left)
-  • Second tree (larger): around x4y13 (just above the first tree)
-However, the actual spatial relationship between grid x2y15 and x4y13 is not consistent with the relative positioning described by the model.
-Figure 10: Hallucination analysis on labeled grids.